### v1

# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to source material remains a significant challenge due to the generation of hallucinations. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucination within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization
[…]
1 Introduction
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of human-like language generation, particularly in tasks like text summarization Chang et al. (2024b). With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Roit et al. (2023); Song et al. (2025). This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al. (2023); Grattafiori et al. (2024); Team et al. (2024); GLM et al. (2024).
Figure 1: Comparison of the faithfulness scores of summaries generated around 50 words, 200 words, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
As LLMs continue to evolve, a paradigm shift is emerging from short-form to long-form generation, enabled by the ability to process extended contexts Wu et al. (2025a). Long-output generation is critical for complex reasoning tasks like long Chain-of-Thought prompting Jaech et al. (2024) and language agents Sumers et al. (2024); Wang et al. (2024), inherently requiring coherent, contextually grounded long responses. While recent benchmarks evaluate long-generation capabilities Wu et al. (2025b); Ye et al. (2025), they often lack contextual grounding, overlooking a crucial aspect of long-form generation: context-aware faithfulness.
A persistent and critical challenge faced by LLMs is the phenomenon of hallucination, wherein the generated content is unfaithful to or unsupported by the input context Huang et al. (2025). While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the spatial distribution of factual errors within the generated sequence. Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al. (2024b); Wan et al. (2025). However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
Motivated by this critical research gap, our work provides the first dedicated investigation into the spatial distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization. Unlike prior work primarily addressing long input contexts or short outputs Tang et al. (2023); Zhang et al. (2024a); Bishop et al. (2024), we tackle long context-aware long response generation, which requires processing lengthy inputs and maintaining faithfulness over a significantly long output sequence. Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness (See Section 3.1 for evaluation metric.) significantly decreases towards the end for long summaries (e.g., 800 words). Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text. To systematically analyze this phenomenon and its implications, our research addresses three core questions:
  1. RQ1.
    Where Do Hallucinations Most Frequently Occur? (§3)
  2. RQ2.
    What Factors Contribute to Hallucination Concentration at the Last Part? (§4)
  3. RQ3.
    How Can We Mitigate the Hallucination in the Last Part? (§5)
To answer these questions, we first empirically characterize the positional distribution of hallucinations (RQ1). Building on this, we explore potential causes of this phenomenon related to the generative process (RQ2). Finally, we propose and evaluate mitigation strategies targeting late-stage hallucinations (RQ3). Our findings highlight the necessity of considering the spatial dimension of errors and provide insights for developing more robust long-form generation systems.
Our contributions can be summarized as follows:
  • We provide the first empirical characterization of the Hallucinate at the Last phenomenon in LLM-based long response generation, particularly in long document summarization.
  • We offer initial insights and analysis into the potential factors contributing to this late-stage hallucination.
  • We investigate mitigation strategies specifically tailored to address positional hallucination.
2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length.
2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 16K tokens. The specific context lengths used in our experiments range from 1K to 8K tokens, with increments of 1K.
2.2 Output Length Categories
We define two categories for output length to study the Hallucinate at the Last phenomenon:
  • Short Output: Summaries with a length between 100 and 200 words.
  • Long Output: Summaries with a length up to 30% of the input context length.
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both short and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
2.3 Overall Faithfulness
To provide a general understanding of the faithfulness levels observed in the generated long summaries before conducting a detailed positional analysis, we first report the overall faithfulness scores.
Context Length Faithfulness(%)
1K 83.3
2K 86.7
3K 78.9
4K 82.9
5K 86.6
6K 85.9
7K 83.9
8K 86.8
Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
As shown in Table 1, we report the overall faithfulness of long summaries generated by Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset, with context lengths ranging from 1K to 8K. We employ the FineSurE Song et al. (2024) framework with GPT-4 Achiam et al. (2023) to compute the overall faithfulness scores. FineSurE evaluates summaries by comparing each sentence with the source context and identifying specific error types when inconsistencies are detected.
As shown in the results, the overall performance is consistent with the results on the CNNDM Hermann et al. (2015) dataset reported in the original FineSurE paper, where the faithfulness of summaries generated by the Llama3-70B-Instruct model is reported to be 85.5%. Building on this, We perform a more fine-grained evaluation of summary faithfulness in Section 3, which includes a positional analysis to examine how factual consistency varies across different parts of the output.
3 Where Do Hallucinations Most Frequently Occur?
In this section, we empirically investigate where hallucinations most frequently occur in long document summarization. We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
Figure 2: Comparison of faithfulness scores for short summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
3.1 Experimental Setup
Models
We conduct experiments using several state-of-the-art Large Language Models to ensure our findings are not model-specific. The models evaluated include Llama3.1-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Albert Q. Jiang et al. (2023), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini Achiam et al. (2023).
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
Datasets
To examine whether the phenomenon is domain-specific, we analyze the spatial distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
3.2 Evaluation Metric
Faithfulness
We adopt an evaluation metric based on atomic facts, which has shown a significant effectiveness in evaluating summary factuality Min et al. (2023); Tang et al. (2024); Scirè et al. (2024); Jing et al. (2024); Wei et al. (2024); Yu et al. (2024).
Let $s_{i}$ be the $i$-th sentence in the generated summary $S$. Each sentence in the generated summary is decomposed into a set of atomic facts $A_{i}$ by utilizing LLM, followed by a filtering process to remove unnecessary atomic facts Yang et al. (2024b), which makes:
$A_{i}=\{a_{i1},a_{i2},...,a_{iNi}\}$ (1)
Using an NLI model, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. We measure only the entailment score, and each atomic fact is assigned the highest entailment score among its comparisons:
$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\textrm{{Entail}}(a_{ij},d_{m})$ (2)
Finally, the average of the entailment scores for a set of filtered atomic facts in a sentence is used as the sentence’s faithfulness score:
$\textrm{Faithfulness}(s_{i})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\textrm{score}(a_{ij})$ (3)
See Appendix B for more experimental details.
Sensitivity
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
3.3 Hallucinate at the Last
The latter part of the summary exhibits the lowest level of faithfulness across nearly all models, and across all contexts and summary lengths.
We present the positional faithfulness discrepancies of both short and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, short summaries generated by the three models—excluding Mistral—consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries, as shown in Figure 3. As the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below 0.6 when 45 sentences are generated from an 8K context.
We report the sensitivity results for long summaries generated by the Llama3.1 model in Table 2, and those for the Mistral model in Table 3. As shown in the results, summaries generated by Llama3.1 tend to exhibit the lowest faithfulness scores in the final bin, resulting in high sensitivity. In contrast, the Mistral model often shows the lowest faithfulness scores in the initial bins, with sensitivity frequently falling below zero. We analyze the underlying causes of these contrasting patterns in Section 4. See Appendix C for more results.
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
Table 2: Comparison of faithfulness scores and Sensitivity for long summaries generated by Llama3.1-8B-Instruct across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
2K 0.75 0.80 0.84 0.82 0.81 -0.8
3K 0.81 0.80 0.82 0.84 0.83 -1.3
4K 0.79 0.81 0.83 0.81 0.81 0.0
5K 0.80 0.84 0.86 0.85 0.79 4.8
6K 0.82 0.87 0.87 0.87 0.86 -0.3
7K 0.82 0.84 0.82 0.83 0.85 -2.3
8K 0.87 0.83 0.87 0.89 0.81 5.5
Table 3: Comparison of faithfulness scores and Sensitivity for long summaries generated by Mistral-7B-Instruct-v0.3 across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, contexts, and summary lengths.
We present the positional faithfulness discrepancy of long summaries generated by the Llama3.1-8B-Instruct model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline. Notably, for the Pubmed dataset, the faithfulness score falls below 0.5 when the model generates 45 sentences at a context length of 8K. Additional sensitivity results across diverse datasets are reported in Appendix D. Notably, across all datasets, none of the computed sensitivity scores fall below zero.
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K. We report results in short summaries in Appendix D.
4 Why Do Hallucinations Frequently Occur at the Last?
This section investigates the underlying factors that contribute to the Hallucinate at the Last phenomenon, exploring two main hypotheses. The first attributes the phenomenon to the inherent nature of summarization: key information is typically concentrated at the beginning of the summary, with less important content appearing toward the end. The second posits that as LLMs generate longer outputs, they increasingly attend to previously generated tokens rather than the original input context, resulting in a shift in attention distribution.
4.1 Is it Intrinsic to Summarization?
This hypothesis is based on the intrinsic structure of human-written and model-generated summaries. In most summarization tasks—especially in news, scientific, and Wikipedia-style documents—the main ideas and most salient information are typically presented early in the summary Hermann et al. (2015); Narayan et al. (2018); Cohan et al. (2018). This structure mirrors the lead bias often observed in source documents, where the beginning contains the most crucial content Kim et al. (2019); Zhao et al. (2022); Ravaut et al. (2024). Motivated by this, our study investigates whether such structural bias also manifests in the form of decreasing factual consistency in generated outputs. Specifically, we compare the positional faithfulness scores of human-written reference summaries and model-generated summaries to assess whether the decline in faithfulness is an artifact of summarization’s inherent structure or a model-specific behavior.
Figure 5: Comparison of faithfulness scores between the reference summary and summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths on the CNNDM dataset.
Experimental Setup
We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens. We compare these references to summaries generated by the Llama3.1-8B-Instruct model, evaluated using the metric described in Section 3.1.
Results & Analysis
Figure 5 presents a positional faithfulness comparison between human-written reference summaries and model-generated summaries from the CNNDM dataset, analyzed across increasing output lengths. As shown in the results, the reference summaries exhibit a different pattern compared to the generated summaries. While the faithfulness scores of the generated summaries drop below 0.6 in the final segment, the reference summaries show a temporary dip around the middle but recover toward the end. This observation suggests that the Hallucination at the Last phenomenon cannot be solely attributed to the inherent characteristics of summarization.
(a) Average attention weights on the first, middle, and last sentences of summaries generated by the Llama3.1-8B-Instruct model.
(b) Average attention weights on the first, middle, and last sentences of summaries generated by the Mistral-7B-Instruct-v0.3 model.
Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 3. The dashed line separates the input context from the generated output.
4.2 Is it Intrinsic to Attention?
Prior research on LLMs has demonstrated that attention weight distributions are closely correlated with the generation process, influencing output coherence Dong et al. (2021); Zhang et al. (2024b). Recent studies on Large Vision-Language Models (LVLMs) further suggest that hallucinations frequently occur in the later parts of generated text Liu et al. (2024a); Lee et al. (2024); Min et al. (2025). This phenomenon has been attributed to a shift in attention: as text generation progresses, attention weights increasingly favor previously generated text tokens over image tokens. Inspired by this observation, we investigate whether a similar trend exists in LLMs by analyzing how attention weights on generated tokens evolve as output length increases.
Experimental Setup
Unlike previous studies that focus on attention to the first generated token Hsieh et al. (2024), we compute attention weights at the sentence level. Moreover, we segment the full sequence into chunks of 100 tokens. Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence. See Appendix E for more details.
Results & Analysis
Figure 6 presents a visualization of average attention weights across sentences. As shown in the results, the Llama3.1-8B-Instruct model—which exhibited the Hallucinate at the Last pattern in Figure 3—assigns nearly four times more attention to the final sentence of the generated summary compared to the first and middle sentences. In contrast, the Mistral-7B-Instruct-v0.3 model assigns similar levels of attention to all three sentence positions. These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination. Moreover, we observe that for most models—excluding Mistral—as output length increases, attention becomes increasingly concentrated on generated tokens, which in turn amplifies hallucination.
5 How to Mitigate the Hallucination at the Last?
We investigate how we can resolve the Hallucinate at the Last phenomenon. To this end, we apply four methods.
Experimental Setup
In this experiment, we generate summaries for the Wikipedia dataset using four different methods, with a context length of 7K tokens. As a baseline, we use summaries generated by the Llama3.1-8B-Instruct model. For comparison, we evaluate the following four methods:
  • BooookScore Chang et al. (2024a) segments the input context into chunks, generates summaries for each chunk individually, and then merges the partial summaries.
  • MInference Jiang et al. (2024) employs sparse attention mechanisms to efficiently process long input sequences.
  • LongWriter-Llama3.1-8B Bai et al. (2025) is a model fine-tuned on a long-output dataset and further enhanced using DPO Rafailov et al. (2023).
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding.
See Appendix F for more experimental details.
Methods Generated Summary Bins Sensitivity
1 2 3 4 5
Llama3.1-8B 0.74 0.75 0.75 0.73 0.64 10.3
+ BooookScore 0.73 0.77 0.78 0.80 0.75 2.0
+ MInference 0.77 0.78 0.78 0.71 0.69 7.0
+ LongWriter 0.85 0.77 0.81 0.63 0.70 6.5
+ AdaCAD 0.79 0.82 0.77 0.81 0.64 15.8
Table 4: Comparison of faithfulness scores and Sensitivity for long summaries generated by different mitigation methods across output bins on the Wikipedia dataset with a 7K context length.
Results & Analysis
We report the faithfulness scores across output bins and the corresponding Sensitivity results in Table 4, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 4, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. However, while BooookScore mitigates the phenomenon, it doesn’t fundamentally solve the problem of generating long responses directly from raw long inputs, relying instead on decomposing the problem via chunking and merging. This highlights the necessity of future work to develop methods that enable models to maintain faithfulness when generating long text directly from long contexts.
6 Related Work
The problem of hallucination in LLMs has been a significant area of research. Prior work has largely focused on developing methods for detecting factual inconsistencies in generated text Chuang et al. (2024); Hu et al. (2024); Kim et al. (2024); Zhong and Litman (2025) and proposing strategies to mitigate their occurrence, often through improved training Zhang et al. (2023); Wan et al. (2023), decoding Shi et al. (2024); Wang et al. (2025), or prompting Zhou et al. (2023). While effective in reducing overall hallucination rates, these studies typically do not analyze the within-sequence distribution of errors.
The concept of positional bias in generation errors has been observed in related domains. Notably, studies on Large Vision-Language Models (LVLMs) have indicated that hallucinations tend to increase towards the end of generated textual descriptions of images Liu et al. (2024a); Lee et al. (2024); Min et al. (2025).
Recently, several studies have begun to explore long-output generation in LLMs Wu et al. (2025b); Ye et al. (2025). However, these works focus on procedural generation tasks that are not grounded in contextual input, limiting their applicability to real-world summarization scenarios. A notable study on multi-document summarization aligns with our findings Belém et al. (2025). Nevertheless, it does not address positional context-aware faithfulness over extended summary lengths.
Our work presents the first dedicated study of where hallucinations occur in LLM-based long document summarization, moving beyond detection to uncover positional patterns, underlying causes, and mitigation strategies.
Figure 7: Comparison of faithfulness scores for long summaries generated by different mitigation methods across increasing output lengths, using the Wikipedia dataset with a 7K context length.
7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into spatially-aware generation techniques.
Limitations
In this study, we explored four domains, primarily because the corresponding datasets provided input contexts of the desired length. However, we were not able to investigate domains such as books, dialogues, movie scripts, or meeting transcripts. Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes. Additionally, we also did not explore models of varying sizes.
Despite these limitations, we believe this work offers an important foundation for future research in the relatively underexplored domain of long-form output generation, particularly in summarization.
Ethics Statement
This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs. All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results. We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary. By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
Appendix A Details for Generating Summaries
Building upon Section 2, this section details our approach to generating long summaries and presents an analysis of the generated outputs.
Prompt
Short outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range. The specific word ranges used for prompting long-summary generation are provided in Table 5, and the prompt template employed for summary generation is shown in Table 7.
Generated Summary Length
Our experiments revealed that controlling the output length of LLMs remains a significant challenge. This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger. In particular, with longer contexts, LLMs frequently generate long summaries that contain repeated sentences or even entire paragraphs. To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
Table 6 presents the average word counts of both short and long summaries generated by each model. As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
Models Avg. Words
short long
Llama3.1-8B-Instruct 279 834
Qwen2.5-7B-Instruct 153 751
Mistral-7B-Instruct-v0.3 285 794
GPT-4o mini 172 925
LongWriter-Llama3.1-8B 398 3291
Table 6: The average word counts of both short and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Evaluation Metric
In this section, we present the experimental details of the evaluation metric used to compute sentence-level faithfulness scores in Section 3.1.
For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al. (2023) [^1] https://github.com/vllm-project/vllm. For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^2] https://huggingface.co/vectara/hallucination_evaluation_model.
Appendix C More Results on Varying Models
Usage of LLMs
We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace. For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18. All summaries were generated using greedy decoding with float16 precision.
More Results
We report the faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset in Table 8, and those of long summaries in Table 9.
Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in short summaries generated by the Llama 3.1 model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Table 8.
We report the faithfulness scores and Sensitivity of short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 10, and those of long summaries in Table 11.
As demonstrated by the experimental results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
Appendix E Attention Weight Calculation
This section provides a more detailed and formal explanation our attention weight computation method introduced in Section 4.2.
Let the full sequence, consisting of the input context and the generated output summary, be defined as:
$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
where $L$ is the number of tokens.
We partition $x_{full}$ into $K$ non-overlapping blocks of 100 tokens each, such that each block is represented as:
$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
for $k\in\{1,2,...,K\}$.
Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
$x^{first}$ =$=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ =$=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ =$=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
We define the average attention weight from block $x^{k}$ to each of the three sentences as follows:
$\begin{gathered}\textrm{{Attn}}_{first}(k)=\frac{1}{100\cdot{T_{1}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{1}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{first})\\ \textrm{{Attn}}_{middle}(k)=\frac{1}{100\cdot{T_{2}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{2}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{middle})\\ \textrm{{Attn}}_{last}(k)=\frac{1}{100\cdot{T_{3}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{3}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{last})\end{gathered}$ (7)
where $\textrm{{attn}}(x_{i}\to x_{j})$ denotes the attention weight from token $x_{i}$ to token $x_{j}$ when generating the output.
These measures quantify the extent to which each 100-token block in the full sequence contributes to the generation of the respective sentences in the output summary.
Appendix F Details for Mitigation Methods
In this section, we provide the experimental details of the mitigation methods described in Section 5.
For the BooookScore method, the input is divided into 2048-token chunks, summaries are generated for each chunk using the Llama3.1-8B-Instruct model, and the partial summaries are then hierarchically merged into a final summary. For LongWriter, we employed the LongWriter-Llama3.1-8B model [^6] https://huggingface.co/THUDM/LongWriter-llama3.1-8b with greedy decoding and bfloat16 precision.
Input Prompt
Write an accurate and engaging summary for the given text in range of {{words_range}} words using only the provided passage (might be irrelevant). Use an unbiased and journalistic tone. Text: {{Text}}
Table 7: Prompt template used to generate summaries from the original document in Section 2.
Figure 8: Comparison of faithfulness scores in short summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Qwen2.5-7B-Instruct 1K 0.78 0.75 0.77 0.72 0.69 6.5
2K 0.83 0.82 0.74 0.72 0.74 3.75
3K 0.82 0.78 0.69 0.81 0.72 5.5
4K 0.85 0.77 0.78 0.72 0.74 4.0
5K 0.83 0.82 0.79 0.80 0.74 7.0
6K 0.79 0.75 0.77 0.75 0.70 6.5
7K 0.80 0.82 0.77 0.80 0.75 4.75
8K 0.84 0.81 0.78 0.82 0.77 4.25
GPT-4o mini 1K 0.79 0.69 0.76 0.73 0.65 9.25
2K 0.79 0.75 0.73 0.66 0.66 7.25
3K 0.84 0.81 0.80 0.73 0.71 8.5
4K 0.74 0.78 0.75 0.72 0.68 6.75
5K 0.86 0.81 0.79 0.75 0.68 12.25
6K 0.77 0.80 0.76 0.77 0.67 10.5
7K 0.87 0.85 0.75 0.77 0.68 13.0
8K 0.80 0.80 0.76 0.77 0.68 10.25
Mistral-7B-Instruct-v0.3 1K 0.83 0.77 0.77 0.75 0.80 -2.0
2K 0.74 0.80 0.80 0.73 0.67 9.75
3K 0.76 0.75 0.78 0.78 0.70 6.75
4K 0.79 0.82 0.75 0.80 0.74 5.0
5K 0.83 0.83 0.84 0.81 0.78 4.75
6K 0.82 0.81 0.82 0.84 0.79 3.25
7K 0.83 0.82 0.88 0.87 0.81 4.0
8K 0.85 0.84 0.83 0.82 0.77 6.5
Table 8: Comparison of faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Qwen2.5-7B-Instruct 1K 0.83 0.76 0.79 0.73 0.70 7.75
2K 0.81 0.79 0.78 0.78 0.70 9.0
3K 0.81 0.78 0.80 0.73 0.70 8.0
4K 0.77 0.76 0.74 0.76 0.69 6.75
5K 0.82 0.81 0.77 0.76 0.71 8.0
6K 0.81 0.78 0.78 0.75 0.67 11.0
7K 0.81 0.74 0.75 0.76 0.75 1.5
8K 0.79 0.78 0.79 0.78 0.73 5.5
GPT-4o mini 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
2K 0.75 0.80 0.84 0.82 0.81 -0.8
3K 0.81 0.80 0.82 0.84 0.83 -1.3
4K 0.79 0.81 0.83 0.81 0.81 0.0
5K 0.80 0.84 0.86 0.85 0.79 4.8
6K 0.82 0.87 0.87 0.87 0.86 -0.3
7K 0.82 0.84 0.82 0.83 0.85 -2.3
8K 0.87 0.83 0.87 0.89 0.81 5.5
Table 9: Comparison of faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.73 0.73 0.75 0.74 0.71 3.0
2K 0.77 0.76 0.75 0.73 0.73 2.25
3K 0.79 0.80 0.76 0.76 0.69 8.75
4K 0.81 0.82 0.81 0.79 0.76 4.75
5K 0.80 0.79 0.79 0.76 0.75 3.5
6K 0.81 0.80 0.80 0.76 0.71 8.25
7K 0.85 0.76 0.78 0.78 0.70 9.25
8K 0.80 0.76 0.73 0.74 0.63 12.75
Arxiv 1K 0.69 0.76 0.79 0.81 0.71 5.25
2K 0.75 0.76 0.80 0.76 0.70 6.75
3K 0.79 0.86 0.84 0.83 0.70 13.0
4K 0.79 0.83 0.83 0.83 0.75 7.0
5K 0.82 0.84 0.84 0.83 0.79 4.25
6K 0.81 0.83 0.81 0.81 0.81 0.5
7K 0.85 0.86 0.86 0.86 0.76 9.75
8K 0.78 0.83 0.83 0.81 0.77 4.25
Pubmed 1K 0.81 0.83 0.81 0.73 0.79 0.5
2K 0.82 0.84 0.78 0.81 0.74 7.25
3K 0.83 0.84 0.81 0.83 0.76 6.75
4K 0.81 0.82 0.78 0.80 0.76 4.25
5K 0.81 0.80 0.83 0.82 0.73 8.5
6K 0.79 0.80 0.86 0.83 0.78 4.0
7K 0.80 0.81 0.82 0.79 0.78 2.5
8K 0.83 0.84 0.81 0.84 0.75 8.0
Govreport 1K 0.72 0.79 0.80 0.82 0.76 2.25
2K 0.83 0.82 0.86 0.83 0.77 6.5
3K 0.83 0.84 0.83 0.82 0.81 2.0
4K 0.82 0.76 0.83 0.78 0.78 1.75
5K 0.83 0.81 0.86 0.82 0.82 1.0
6K 0.84 0.83 0.87 0.86 0.82 3.0
7K 0.83 0.86 0.89 0.89 0.79 7.75
8K 0.84 0.82 0.87 0.85 0.81 3.5
Table 10: Comparison of faithfulness scores and Sensitivity for short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
Datasets Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Wikipedia 1K 0.76 0.75 0.69 0.68 0.68 4.0
2K 0.73 0.76 0.73 0.70 0.63 10.0
3K 0.71 0.72 0.72 0.74 0.62 10.25
4K 0.71 0.75 0.70 0.70 0.67 4.5
5K 0.71 0.75 0.79 0.79 0.70 6.0
6K 0.69 0.76 0.73 0.76 0.61 12.5
7K 0.78 0.79 0.75 0.76 0.67 10.0
8K 0.74 0.79 0.74 0.77 0.67 9.0
Arxiv 1K 0.77 0.82 0.81 0.80 0.74 6.0
2K 0.76 0.78 0.80 0.72 0.71 5.5
3K 0.75 0.79 0.81 0.78 0.68 10.25
4K 0.71 0.81 0.74 0.77 0.73 2.75
5K 0.77 0.83 0.85 0.81 0.74 7.5
6K 0.69 0.82 0.79 0.75 0.75 1.25
7K 0.86 0.86 0.80 0.80 0.77 6.0
8K 0.79 0.81 0.85 0.82 0.73 8.75
Pubmed 1K 0.75 0.85 0.84 0.80 0.78 3.0
2K 0.76 0.83 0.82 0.83 0.72 9.0
3K 0.81 0.85 0.86 0.78 0.76 6.5
4K 0.77 0.82 0.79 0.79 0.68 11.25
5K 0.78 0.81 0.81 0.78 0.70 9.5
6K 0.80 0.87 0.83 0.77 0.74 7.75
7K 0.85 0.85 0.84 0.79 0.77 6.25
8K 0.80 0.79 0.81 0.77 0.67 12.25
Govreport 1K 0.80 0.80 0.85 0.87 0.78 5.0
2K 0.82 0.85 0.83 0.85 0.78 5.75
3K 0.81 0.86 0.89 0.81 0.82 2.25
4K 0.76 0.86 0.86 0.82 0.77 5.5
5K 0.82 0.85 0.88 0.82 0.78 6.25
6K 0.86 0.88 0.88 0.87 0.80 7.25
7K 0.86 0.91 0.89 0.88 0.83 5.5
8K 0.84 0.87 0.89 0.84 0.80 6.0
Table 11: Comparison of faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K. The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.

### v2

@@
 Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs.
-However, faithfulness to source material remains a significant challenge due to the generation of hallucinations.
-While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucination within generated text, particularly in long outputs.
+However, faithfulness to the source material remains a significant challenge due to the generation of hallucinations by LLMs.
+While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs.
 In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study.
@@
 Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization
-[…]
+Joonho Yang1, Seunghyun Yoon2, Hwan Chang1, Byeongjeong Kim1, Hwanhee Lee1thanks: Corresponding author.
+1Department of Artificial Intelligence, Chung-Ang University, 2Adobe Research, USA {plm3332, hwanchang, michael97k, hwanheelee}@cau.ac.kr, syoon@adobe.com
 1 Introduction
@@
 (2024b).
-With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Roit et al.
-(2023); Song et al.
+With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Song et al.
 (2025).
 This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al.
-(2023); Grattafiori et al.
 (2023); Team et al.
-(2024); GLM et al.
 (2024).
-Figure 1: Comparison of the faithfulness scores of summaries generated around 50 words, 200 words, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
+Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
 As LLMs continue to evolve, a paradigm shift is emerging from short-form to long-form generation, enabled by the ability to process extended contexts Wu et al.
@@
 (2025).
-While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the spatial distribution of factual errors within the generated sequence.
+While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the positional distribution of factual errors within the generated sequence.
 Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al.
@@
 However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
-Motivated by this critical research gap, our work provides the first dedicated investigation into the spatial distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization.
-Unlike prior work primarily addressing long input contexts or short outputs Tang et al.
-(2023); Zhang et al.
+Motivated by this critical research gap, our work provides the first dedicated investigation into the positional distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization.
+Unlike prior work primarily addressing long input contexts or short outputs Zhang et al.
 (2024a); Bishop et al.
@@
 Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
-Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness (See Section 3.1 for evaluation metric.) significantly decreases towards the end for long summaries (e.g., 800 words).
+Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness significantly decreases towards the end for long summaries (e.g., 800 words).
 Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text.
@@
     How Can We Mitigate the Hallucination in the Last Part? (§5)
-To answer these questions, we first empirically characterize the positional distribution of hallucinations (RQ1).
-Building on this, we explore potential causes of this phenomenon related to the generative process (RQ2).
-Finally, we propose and evaluate mitigation strategies targeting late-stage hallucinations (RQ3).
-Our findings highlight the necessity of considering the spatial dimension of errors and provide insights for developing more robust long-form generation systems.
+To answer these questions, we empirically characterize the positional distribution of hallucinations (RQ1), investigate their generative causes (RQ2), and propose mitigation strategies (RQ3).
+Our results underscore the importance of the positional dimension of errors in building robust long-form generation systems.
 Our contributions can be summarized as follows:
@@
 We prompt LLMs to generate summaries from original documents, varying both the context length and the output length.
+For all experiments, we generate summaries using greedy decoding.
+The prompt is provided in Appendix A.
 2.1 Input Context Lengths
-To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 16K tokens.
-The specific context lengths used in our experiments range from 1K to 8K tokens, with increments of 1K.
+To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 8K tokens.
+For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
 2.2 Output Length Categories
 We define two categories for output length to study the Hallucinate at the Last phenomenon:
-  • Short Output: Summaries with a length between 100 and 200 words.
-  • Long Output: Summaries with a length up to 30% of the input context length.
+  • Standard Summary: Summaries with a length between 100 and 200 words.
+[^1] The length of the standard summaries in our study was determined based on the average length of the reference summaries in the dataset, which is employed in Section 4.
+  • Long Summary: Summaries with a length up to 30% of the input context length.
 This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses.
-By including both short and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases.
+By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases.
 See Appendix A for more details.
 2.3 Overall Faithfulness
-To provide a general understanding of the faithfulness levels observed in the generated long summaries before conducting a detailed positional analysis, we first report the overall faithfulness scores.
+As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al.
+(2024) to ensure consistent quality across different context lengths, prior to our main positional analysis.
+Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al.
+(2024) on the Wikipedia dataset (1K–8K context lengths), evaluated using FineSurE with GPT-4 Achiam et al.
+(2023).
+FineSurE assesses each sentence against the source, identifying error types when inconsistencies occur.
+The results align with those on CNNDM Hermann et al.
+(2015), where Llama3-70B-Instruct achieved 85.5% faithfulness.
+We next conduct a finer-grained positional analysis in Section 3 for various setups.
 Context Length Faithfulness(%)
@@
 Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
-As shown in Table 1, we report the overall faithfulness of long summaries generated by Llama3.1-8B-Instruct Grattafiori et al.
-(2024) on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
-We employ the FineSurE Song et al.
-(2024) framework with GPT-4 Achiam et al.
-(2023) to compute the overall faithfulness scores.
-FineSurE evaluates summaries by comparing each sentence with the source context and identifying specific error types when inconsistencies are detected.
-As shown in the results, the overall performance is consistent with the results on the CNNDM Hermann et al.
-(2015) dataset reported in the original FineSurE paper, where the faithfulness of summaries generated by the Llama3-70B-Instruct model is reported to be 85.5%.
-Building on this, We perform a more fine-grained evaluation of summary faithfulness in Section 3, which includes a positional analysis to examine how factual consistency varies across different parts of the output.
 3 Where Do Hallucinations Most Frequently Occur?
@@
 We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
-Figure 2: Comparison of faithfulness scores for short summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
+Figure 2: Comparison of faithfulness scores for standard summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
 3.1 Experimental Setup
 Models
-We conduct experiments using several state-of-the-art Large Language Models to ensure our findings are not model-specific.
-The models evaluated include Llama3.1-8B-Instruct Grattafiori et al.
-(2024), Mistral-7B-Instruct-v0.3 Albert Q.
-Jiang et al.
+We evaluate our approach on multiple state-of-the-art LLMs, including Llama3.1-8B-Instruct, Gemma3-12B-Instruct Team et al.
 (2025), Qwen2.5-7B-Instruct Yang et al.
-(2024a), and GPT-4o mini Achiam et al.
-(2023).
-Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
+(2024a), and GPT-4o mini, to ensure model-agnostic findings.
+For Gemma, we retain the default sliding window attention context length of 1024 tokens, while for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
 Datasets
-To examine whether the phenomenon is domain-specific, we analyze the spatial distribution of faithfulness within generated summaries across various domains.
+To examine whether the phenomenon is domain-specific, we analyze the positional distribution of faithfulness within generated summaries across various domains.
 In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al.
@@
 Faithfulness
-We adopt an evaluation metric based on atomic facts, which has shown a significant effectiveness in evaluating summary factuality Min et al.
-(2023); Tang et al.
-(2024); Scirè et al.
+To enable an in-depth analysis of positional bias within each generated summary, we use the MiniCheck-Flan-T5-Large Tang et al.
+(2024) model to evaluate faithfulness.
+MiniCheck is particularly suitable for this granular task, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al.
 (2023); Jing et al.
-(2024); Wei et al.
 (2024); Yu et al.
@@
 (2024b), which makes:
-$A_{i}=\{a_{i1},a_{i2},...,a_{iNi}\}$ (1)
-Using an NLI model, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner.
-We measure only the entailment score, and each atomic fact is assigned the highest entailment score among its comparisons:
-$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\textrm{{Entail}}(a_{ij},d_{m})$ (2)
-Finally, the average of the entailment scores for a set of filtered atomic facts in a sentence is used as the sentence’s faithfulness score:
+$A_{i}=\{a_{i1},a_{i2},...,a_{iN_{i}}\}$ (1)
+Using a MiniCheck model, which is denoted as $\mathcal{M}$, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner.
+Each atomic fact is assigned the highest score among its comparisons:
+$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\mathcal{M}(d_{m},a_{ij})$ (2)
+Finally, we use the average of the scores for a set of filtered atomic facts in a sentence as its faithfulness score:
 $\textrm{Faithfulness}(s_{i})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\textrm{score}(a_{ij})$ (3)
-See Appendix B for more experimental details.
+Additional experimental details and an example of the evaluation procedure are provided in Appendix B.
+Furthermore, we verify the quality of atomic facts generated by the LLM via human evaluation in Table 2 for atomic facts extracted from long summaries generated by Llama.
+Context Length Validation(%)
+1K 98.55
+2K 99.08
+3K 97.62
+4K 95.92
+5K 99.01
+6K 96.18
+7K 99.12
+8K 99.64
+Table 2: Human evaluation of the quality (completeness) of atomic facts.
+For each context length, we randomly sampled 10 summaries and conducted the evaluation manually.
+Each atomic fact was annotated as either true or false depending on whether it is present in the summary.
 Sensitivity
 In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs.
-Specifically, each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed.
+Specifically, following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins or segments, and the average faithfulness score of the atomic facts within each bin is computed.
 We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins.
@@
 The larger the sensitivity, the more pronounced the hallucination at the end of the output.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct Overall Avg. 0.81 0.82 0.79 0.80 0.75 5.5
+12K 0.80 0.79 0.66 0.63 0.69 9.5
+16K 0.86 0.83 0.85 0.83 0.79 5.25
+Llama3.1-70B-Instruct Overall Avg. 0.83 0.82 0.83 0.81 0.77 5.25
+Gemma3-4B-Instruct Overall Avg. 0.85 0.82 0.83 0.85 0.79 4.75
+Gemma3-12B-Instruct Overall Avg. 0.85 0.83 0.84 0.83 0.79 4.75
+GPT-4o mini Overall Avg. 0.83 0.81 0.81 0.78 0.76 4.75
+Qwen2.5-7B-Instruct Overall Avg. 0.86 0.83 0.83 0.82 0.84 -0.5
+Table 3: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset.
+Overall Avg.
+denotes the average faithfulness score aggregated across context lengths from 1K to 8K.
 3.3 Hallucinate at the Last
-The latter part of the summary exhibits the lowest level of faithfulness across nearly all models, and across all contexts and summary lengths.
-We present the positional faithfulness discrepancies of both short and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively.
-As shown in Figure 2, short summaries generated by the three models—excluding Mistral—consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end.
+Model Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.92 0.90 0.91 0.87 0.68 22.00
+2K 0.82 0.74 0.83 0.72 0.71 6.75
+3K 0.94 0.94 0.81 0.86 0.68 20.75
+4K 0.92 0.92 0.88 0.87 0.82 7.75
+5K 0.91 0.88 0.89 0.78 0.70 16.50
+6K 0.86 0.84 0.79 0.89 0.73 11.50
+7K 0.97 0.89 0.82 0.94 0.66 24.50
+8K 0.93 0.90 0.86 0.79 0.65 22.00
+Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset.
+Each atomic fact was annotated as either true or false depending on whether it is present in the document.
+The results were manually evaluated and aggregated.
+The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across various contexts and summary lengths.
+The decline becomes most pronounced when the context length reaches 8K and the summaries are long.
+We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively.
+As shown in Figure 2, standard summaries generated by the three models, excluding Qwen, consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end.
 Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length.
 Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
-However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries, as shown in Figure 3.
-As the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below 0.6 when 45 sentences are generated from an 8K context.
-We report the sensitivity results for long summaries generated by the Llama3.1 model in Table 2, and those for the Mistral model in Table 3.
-As shown in the results, summaries generated by Llama3.1 tend to exhibit the lowest faithfulness scores in the final bin, resulting in high sensitivity.
-In contrast, the Mistral model often shows the lowest faithfulness scores in the initial bins, with sensitivity frequently falling below zero.
-We analyze the underlying causes of these contrasting patterns in Section 4.
-See Appendix C for more results.
-Model Context Generated Summary Bins Sensitivity
-Length 1 2 3 4 5
-Llama3.1-8B-Instruct 1K 0.77 0.74 0.69 0.69 0.69 3.3
-2K 0.74 0.72 0.70 0.66 0.63 7.5
-3K 0.75 0.70 0.69 0.65 0.58 11.8
-4K 0.74 0.71 0.68 0.69 0.64 6.5
-5K 0.73 0.72 0.72 0.68 0.60 11.3
-6K 0.74 0.72 0.66 0.65 0.61 8.3
-7K 0.73 0.70 0.69 0.66 0.62 7.5
-8K 0.76 0.70 0.72 0.69 0.61 10.8
-Table 2: Comparison of faithfulness scores and Sensitivity for long summaries generated by Llama3.1-8B-Instruct across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
-Model Context Generated Summary Bins Sensitivity
-Length 1 2 3 4 5
-Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
-2K 0.75 0.80 0.84 0.82 0.81 -0.8
-3K 0.81 0.80 0.82 0.84 0.83 -1.3
-4K 0.79 0.81 0.83 0.81 0.81 0.0
-5K 0.80 0.84 0.86 0.85 0.79 4.8
-6K 0.82 0.87 0.87 0.87 0.86 -0.3
-7K 0.82 0.84 0.82 0.83 0.85 -2.3
-8K 0.87 0.83 0.87 0.89 0.81 5.5
-Table 3: Comparison of faithfulness scores and Sensitivity for long summaries generated by Mistral-7B-Instruct-v0.3 across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
-The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, contexts, and summary lengths.
-We present the positional faithfulness discrepancy of long summaries generated by the Llama3.1-8B-Instruct model across different datasets in Figure 4.
+Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries except for Qwen, as shown in Figure 3.
+For these models, as the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below a faithfulness score of 0.70 when 45 sentences are generated from an 8K context.
+Qwen, in contrast, often exhibits the largest drop in faithfulness in the middle of the summary, after which the scores recover toward the end.
+Tables 3 report the faithfulness and sensitivity of various models.
+We aggregate the faithfulness scores across context lengths from 1K to 8K and report the overall average.
+All models except Qwen show markedly low faithfulness scores in the final bin, reflecting high sensitivity.
+Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size or context length.
+Qwen, however, is the only model that records sensitivity values below zero.
+Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity, and we provide a detailed analysis of these contrasting patterns in Section 4.
+Supplementary results across context lengths and methods are provided in Appendix C.
+The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, with a marked decline once the context length exceeds 7K.
+We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 4.
 As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains.
 Faithfulness scores tend to decrease toward the final segments as the output length increases.
-In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline.
-Notably, for the Pubmed dataset, the faithfulness score falls below 0.5 when the model generates 45 sentences at a context length of 8K.
-Additional sensitivity results across diverse datasets are reported in Appendix D.
-Notably, across all datasets, none of the computed sensitivity scores fall below zero.
+In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline, with the deterioration becoming especially pronounced once the context length exceeds 7K.
+Additional sensitivity results across diverse datasets are provided in Appendix D.
 Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
-We report results in short summaries in Appendix D.
 4 Why Do Hallucinations Frequently Occur at the Last?
@@
 4.1 Is it Intrinsic to Summarization?
-This hypothesis is based on the intrinsic structure of human-written and model-generated summaries.
-In most summarization tasks—especially in news, scientific, and Wikipedia-style documents—the main ideas and most salient information are typically presented early in the summary Hermann et al.
+This hypothesis draws on the intrinsic structure of both human- and model-generated summaries.
+In typical summarization domains such as news, scientific articles, and Wikipedia-style documents, key information is predominantly introduced early in the summary Hermann et al.
 (2015); Narayan et al.
 (2018); Cohan et al.
-(2018).
-This structure mirrors the lead bias often observed in source documents, where the beginning contains the most crucial content Kim et al.
+(2018), reflecting the lead bias observed in source texts, where the opening contains the most salient content Kim et al.
 (2019); Zhao et al.
@@
 (2024).
-Motivated by this, our study investigates whether such structural bias also manifests in the form of decreasing factual consistency in generated outputs.
-Specifically, we compare the positional faithfulness scores of human-written reference summaries and model-generated summaries to assess whether the decline in faithfulness is an artifact of summarization’s inherent structure or a model-specific behavior.
-Figure 5: Comparison of faithfulness scores between the reference summary and summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths on the CNNDM dataset.
+Building on this, our study investigates whether this structural bias is also reflected in decreasing factual consistency within generated outputs by comparing positional faithfulness scores between human-written and model-generated summaries.
+Figure 5: Faithfulness scores of reference summaries and those generated by Llama on the CNNDM dataset.
 Experimental Setup
 We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens.
-We compare these references to summaries generated by the Llama3.1-8B-Instruct model, evaluated using the metric described in Section 3.1.
+We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
 Results & Analysis
-Figure 5 presents a positional faithfulness comparison between human-written reference summaries and model-generated summaries from the CNNDM dataset, analyzed across increasing output lengths.
-As shown in the results, the reference summaries exhibit a different pattern compared to the generated summaries.
-While the faithfulness scores of the generated summaries drop below 0.6 in the final segment, the reference summaries show a temporary dip around the middle but recover toward the end.
-This observation suggests that the Hallucination at the Last phenomenon cannot be solely attributed to the inherent characteristics of summarization.
-(a) Average attention weights on the first, middle, and last sentences of summaries generated by the Llama3.1-8B-Instruct model.
-(b) Average attention weights on the first, middle, and last sentences of summaries generated by the Mistral-7B-Instruct-v0.3 model.
-Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 3. The dashed line separates the input context from the generated output.
+Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries.
+We observe that the faithfulness of the generated summaries declines steadily, falling below 0.6 in the final segment, whereas reference summaries dip around the middle but recover.
+This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
 4.2 Is it Intrinsic to Attention?
@@
 Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence.
-See Appendix E for more details.
+Details of the attention weight calculation are provided in Appendix E.
 Results & Analysis
 Figure 6 presents a visualization of average attention weights across sentences.
-As shown in the results, the Llama3.1-8B-Instruct model—which exhibited the Hallucinate at the Last pattern in Figure 3—assigns nearly four times more attention to the final sentence of the generated summary compared to the first and middle sentences.
-In contrast, the Mistral-7B-Instruct-v0.3 model assigns similar levels of attention to all three sentence positions.
+As shown in the results, the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 3, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences.
+In contrast, the Qwen model assigns similar levels of attention to all three sentence positions.
 These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
-Moreover, we observe that for most models—excluding Mistral—as output length increases, attention becomes increasingly concentrated on generated tokens, which in turn amplifies hallucination.
-5 How to Mitigate the Hallucination at the Last?
-We investigate how we can resolve the Hallucinate at the Last phenomenon. To this end, we apply four methods.
+This consistent distribution in Qwen can be attributed to its sliding window attention mechanism Child et al.
+(2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations.
+Table 5 compares the Qwen model with and without sliding window attention.
+In this setting, the sliding window context length is fixed at 16000 tokens.
+The results clearly show that sliding window attention omission increases sensitivity across all context and summary lengths, up to 9.75.
+They demonstrate that employing sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
+(a) Average attention weights of summaries generated by Llama.
+(b) Average attention weights of summaries generated by Qwen.
+Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 3. The dashed line separates the input context from the generated output.
+5 How to Mitigate the Hallucinate at the Last?
+In the preceding section, we demonstrated that the Hallucinate at the Last phenomenon could be partially mitigated through the use of sliding window attention.
+In the following analysis, we investigate strategies to address this issue in models that are not trained with sliding window attention.
+To this end, we apply four methods that may mitigate the phenomenon.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Standard Summary
+Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
+2K 0.90 0.89 0.83 0.78 0.84 1.0
+3K 0.87 0.84 0.79 0.83 0.82 1.25
+4K 0.91 0.88 0.90 0.79 0.89 -2.0
+5K 0.86 0.89 0.85 0.76 0.89 -5.0
+6K 0.93 0.91 0.92 0.80 0.90 -1.0
+7K 0.90 0.89 0.89 0.84 0.82 6.0
+8K 0.89 0.93 0.88 0.89 0.92 -2.25
+Without 1K 0.88 0.79 0.74 0.76 0.84 -4.75(+1.5)
+2K 0.90 0.89 0.83 0.78 0.84 1.0(+0.0)
+Sliding 3K 0.88 0.86 0.82 0.85 0.83 2.25(+1.0)
+4K 0.91 0.88 0.87 0.83 0.83 4.25(+6.25)
+Winodw 5K 0.87 0.87 0.81 0.82 0.80 4.25(+9.25)
+6K 0.93 0.90 0.91 0.85 0.86 3.75(+4.75)
+Attention 7K 0.91 0.91 0.89 0.84 0.79 9.75(+3.75)
+8K 0.88 0.93 0.89 0.89 0.92 -2.25(+0.0)
+Long Summary
+Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
+2K 0.83 0.89 0.84 0.83 0.82 2.75
+3K 0.83 0.84 0.86 0.79 0.84 -1.0
+4K 0.87 0.83 0.79 0.83 0.84 -1.0
+5K 0.87 0.89 0.89 0.80 0.85 1.25
+6K 0.89 0.81 0.78 0.78 0.82 -0.5
+7K 0.86 0.78 0.79 0.83 0.84 -2.5
+8K 0.86 0.85 0.85 0.85 0.85 0.25
+Without 1K 0.89 0.77 0.83 0.84 0.80 3.25(+4.25)
+2K 0.83 0.89 0.83 0.80 0.78 5.75(+3.0)
+Sliding 3K 0.83 0.84 0.86 0.79 0.84 -1.0(+0.0)
+4K 0.87 0.83 0.79 0.83 0.84 -1.0(+0.0)
+Winodw 5K 0.86 0.90 0.83 0.79 0.80 4.5(+3.25)
+6K 0.89 0.82 0.84 0.83 0.78 6.5(+7.0)
+Attention 7K 0.86 0.80 0.78 0.83 0.82 -0.25(+2.25)
+8K 0.88 0.85 0.85 0.86 0.82 4.0(+3.75)
+Table 5: Faithfulness scores and sensitivity for standard and long summaries generated by Qwen2.5-7B-Instruct on the Wikipedia dataset, with and without sliding window attention.
+Note that higher sensitivity indicates stronger hallucination effects toward the end of the output.
 Experimental Setup
-In this experiment, we generate summaries for the Wikipedia dataset using four different methods, with a context length of 7K tokens.
-As a baseline, we use summaries generated by the Llama3.1-8B-Instruct model.
-For comparison, we evaluate the following four methods:
+In this experiment, we generate Wikipedia summaries with a 7K context using four methods, taking Llama3.1-8B-Instruct as the baseline.
+For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
   • BooookScore Chang et al. (2024a) segments the input context into chunks, generates summaries for each chunk individually, and then merges the partial summaries.
@@
   • LongWriter-Llama3.1-8B Bai et al. (2025) is a model fine-tuned on a long-output dataset and further enhanced using DPO Rafailov et al. (2023).
-  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding.
-See Appendix F for more experimental details.
+  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024).
 Methods Generated Summary Bins Sensitivity
@@
 + AdaCAD 0.79 0.82 0.77 0.81 0.64 15.8
-Table 4: Comparison of faithfulness scores and Sensitivity for long summaries generated by different mitigation methods across output bins on the Wikipedia dataset with a 7K context length.
+Table 6: Faithfulness scores and sensitivity for long summaries generated by different mitigation methods on the Wikipedia dataset.
 Results & Analysis
-We report the faithfulness scores across output bins and the corresponding Sensitivity results in Table 4, and the faithfulness scores across increasing output lengths for each method in Figure 7.
+We report the faithfulness scores across output bins and the sensitivity results in Table 6, and the faithfulness scores across increasing output lengths for each method in Figure 7.
 As shown in Table 6, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods.
@@
 These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon.
-However, while BooookScore mitigates the phenomenon, it doesn’t fundamentally solve the problem of generating long responses directly from raw long inputs, relying instead on decomposing the problem via chunking and merging.
-This highlights the necessity of future work to develop methods that enable models to maintain faithfulness when generating long text directly from long contexts.
+Nevertheless, while BooookScore serves as a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation.
+This highlights the need for future work to develop methods that maintain faithfulness when generating long text directly from long contexts.
 6 Related Work
@@
 A notable study on multi-document summarization aligns with our findings Belém et al.
+(2025); Zhao et al.
 (2025).
@@
 Our work presents the first dedicated study of where hallucinations occur in LLM-based long document summarization, moving beyond detection to uncover positional patterns, underlying causes, and mitigation strategies.
-Figure 7: Comparison of faithfulness scores for long summaries generated by different mitigation methods across increasing output lengths, using the Wikipedia dataset with a 7K context length.
+Figure 7: Faithfulness scores for long summaries generated by different mitigation methods on the Wikipedia dataset across increasing output lengths.
 7 Conclusion
@@
 We investigated the contributing factors and explored targeted mitigation strategies.
-Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into spatially-aware generation techniques.
+Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into positionally-aware generation techniques.
 Limitations
@@
 Books were excluded due to their excessive length, while the other domains lacked datasets with suitable input lengths or sufficient sample sizes.
-Additionally, we also did not explore models of varying sizes.
 Despite these limitations, we believe this work offers an important foundation for future research in the relatively underexplored domain of long-form output generation, particularly in summarization.
-Ethics Statement
-This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs.
-All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results.
-We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary.
-By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
 Acknowledgement
-This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligent Graduate School Program (Chung-Ang University)].
+This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
 Appendix A Details for Generating Summaries
@@
 Prompt
-Short outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range.
+Standard outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range.
 The specific word ranges used for prompting long-summary generation are provided in Table 7, and the prompt template employed for summary generation is shown in Table 10.
+Context Length Words Range
+1K 200 to 250
+2K 400 to 500
+3K 600 to 750
+4K 800 to 1000
+5K 1000 to 1250
+6K 1200 to 1500
+7K 1400 to 1750
+8K 1600 to 2000
+Table 7: Words range used in the prompt for long output generation.
 Generated Summary Length
@@
 To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
-Table 6 presents the average word counts of both short and long summaries generated by each model.
+Table 8 presents the average word counts of both standard and long summaries generated by each model.
 As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
 Models Avg. Words
-short long
-Llama3.1-8B-Instruct 279 834
-Qwen2.5-7B-Instruct 153 751
-Mistral-7B-Instruct-v0.3 285 794
-GPT-4o mini 172 925
-LongWriter-Llama3.1-8B 398 3291
-Table 6: The average word counts of both short and long summaries generated by each model on the Wikipedia 7K dataset.
-Appendix B Details for Evaluation Metric
-In this section, we present the experimental details of the evaluation metric used to compute sentence-level faithfulness scores in Section 3.1.
+standard long
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 8: The average word counts of both standard and long summaries generated by each model on the Wikipedia 7K dataset.
+Appendix B Details for Model Usage
+In this section, we present the experimental details for the model usage in Section 3.
+Details for Evaluation Method
 For atomic fact decomposition, we utilize the Llama3.1-8B-Instruct model with vLLM Kwon et al.
 (2023) [^2] https://github.com/vllm-project/vllm.
-For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al.
-(2024) [^2] https://huggingface.co/vectara/hallucination_evaluation_model.
-Appendix C More Results on Varying Models
-Usage of LLMs
-We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace.
+To further illustrate the evaluation process, we present in Table 9 a concrete example that demonstrates how a given sentence is decomposed into its constituent atomic facts and subsequently assessed for faithfulness.
+This example clarifies the step-by-step procedure and provides transparency regarding the alignment between the generated atomic facts and the underlying reference information.
+First Sentence
+Sara Sorribes Tormo is a talented Spanish professional tennis player born on October 8, 1996, in Castellón de la Plana, Spain.
+Atomic Facts Score
+Sara Sorribes Tormo is a Spanish professional tennis player. 0.92
+Sara Sorribes Tormo was born in Spain. 0.90
+Sara Sorribes Tormo was born on October 8. 0.87
+Sara Sorribes Tormo was born in 1996. 0.96
+Sara Sorribes Tormo was born in Castellón de la Plana. 0.93
+Avg. Score 0.92
+Last Sentence
+As Sara navigates her career, her unique blend of skill, perseverance, and spirit positions her as a promising athlete, inspiring both current and aspiring players.
+Atomic Facts Score
+Sara navigates her career. 0.87
+Sara’s career is that of an athlete. 0.91
+Sara has a unique blend of skill. 0.14
+Sara has perseverance. 0.81
+Sara is a promising athlete. 0.30
+Sara inspires current players. 0.11
+Sara inspires aspiring players. 0.15
+Avg. Score 0.47
+Table 9: Example of the evaluation procedure with the MiniCheck-Flan-T5-Large Model.
+Details for LLMs Usage
+We utilized publicly available instruction-tuned models, including Llama3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Gemma3-12B-Instruct [^4] https://huggingface.co/google/gemma-3-12b-it, and Qwen2.5-7B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct from HuggingFace.
 For the GPT-4o mini model, we used gpt-4o-mini-2024-07-18.
 All summaries were generated using greedy decoding with float16 precision.
-More Results
-We report the faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset in Table 8, and those of long summaries in Table 9.
+Appendix C More Results with Different Methods
+We adopt a diverse set of methods to validate the robustness of our faithfulness evaluation framework.
+In this section, we present experimental results for faithfulness assessment models beyond MiniCheck, incorporating approaches based on NLI, ROUGE-L, and LLMs.
+Specifically, we substitute the scoring model $\mathcal{M}$ in Equation 2 of Section 3 with these alternative methods.
+Minicheck Based
+We utilize MiniCheck-Flan-T5-Large Model [^6] https://huggingface.co/lytang/MiniCheck-Flan-T5-Large for the main evaluation method.
+We report the faithfulness scores and sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 11, and those of long summaries in Table 12.
+NLI Based
+For the NLI model, we use the state-of-the-art hallucination evaluation model Bao et al. (2024) [^7] https://huggingface.co/vectara/hallucination_evaluation_model.
+We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 13, and those of long summaries in Table 14.
+As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit high sensitivity, whereas the Qwen model occasionally demonstrates sensitivity values even below zero.
+ROUGE Based
+We report the faithfulness scores and sensitivity assessed by ROUGE-L Lin (2004) for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 17, and those of long summaries in Table 18.
+LLM Based
+We present the error rates assessed by FineSurE framework for standard summaries generated by different models across output bins on the Wikipedia dataset in Table 15, and those of long summaries in Table 16.
+As observed in the experimental results, consistent with the main results, the Llama and GPT models consistently exhibit the highest error rates in the final bin.
+In contrast, the Qwen model does not yield the highest error rates in the final bin for both the cases of standard and long summaries.
 Appendix D More Results on Varying Datasets
-We report comparison of faithfulness scores in short summaries generated by the Llama 3.1 model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Table 8.
-We report the faithfulness scores and Sensitivity of short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 10, and those of long summaries in Table 11.
-As demonstrated by the experimental results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
+We report comparison of faithfulness scores in standard summaries generated by the Llama model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Figure 8.
+Minicheck Based
+We report the faithfulness scores and sensitivity assessed by MiniCheck-Flan-T5-Large model of standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 19, and those of long summaries in Table 20.
+The experimental results reveal that in the case of standard summaries, sensitivities below zero are frequently observed; notably, in the Pubmed dataset, more than half of the values fall below zero.
+In contrast, such negative sensitivities are rarely observed in long summaries, with the Pubmed dataset in particular showing no sensitivities below zero, unlike the standard summary setting.
+NLI Based
+We report the faithfulness scores and sensitivity assessed by NLI model for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains in Table 21, and those of long summaries in Table 22.
+As demonstrated by the experimental results, consistent with the main results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
 Appendix E Attention Weight Calculation
-This section provides a more detailed and formal explanation our attention weight computation method introduced in Section 4.2.
+This section provides a more detailed and formal explanation of our attention weight computation method introduced in Section 4.
 Let the full sequence, consisting of the input context and the generated output summary, be defined as:
@@
 Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
-$x^{first}$ =$=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
-$x^{second}$ =$=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
-$x^{third}$ =$=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
+$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
+$x^{second}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
+$x^{third}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
 where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
@@
 For LongWriter, we employed the LongWriter-Llama3.1-8B model [^8] https://huggingface.co/THUDM/LongWriter-llama3.1-8b with greedy decoding and bfloat16 precision.
+Appendix G Use of Large Language Model
+This paper primarily utilized the AI assistant (ChatGPT-5) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
 Input Prompt
@@
 Table 10: Prompt template used to generate summaries from the original document in Section 2.
-Figure 8: Comparison of faithfulness scores in short summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
+Figure 8: Comparison of faithfulness scores in standard summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.81 0.79 0.85 0.82 0.79 2.75
+2K 0.84 0.79 0.81 0.81 0.77 4.25
+3K 0.87 0.85 0.84 0.83 0.79 5.75
+4K 0.91 0.90 0.88 0.86 0.88 0.75
+5K 0.85 0.85 0.86 0.83 0.83 1.75
+6K 0.89 0.85 0.85 0.83 0.81 4.5
+7K 0.91 0.81 0.85 0.87 0.81 5.0
+8K 0.87 0.84 0.82 0.82 0.79 4.75
+Llama3.1-70B-Instruct 1K 0.87 0.83 0.86 0.80 0.77 7.0
+2K 0.92 0.81 0.83 0.83 0.72 12.75
+3K 0.90 0.82 0.86 0.77 0.83 0.75
+4K 0.94 0.85 0.85 0.86 0.77 10.5
+5K 0.91 0.86 0.85 0.83 0.89 -2.75
+6K 0.89 0.85 0.84 0.87 0.83 3.25
+7K 0.87 0.88 0.90 0.83 0.84 3.0
+8K 0.92 0.91 0.84 0.84 0.85 2.75
+Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
+2K 0.90 0.89 0.83 0.78 0.84 1.0
+3K 0.87 0.84 0.79 0.83 0.82 1.25
+4K 0.91 0.88 0.90 0.79 0.89 -2.0
+5K 0.86 0.89 0.85 0.76 0.89 -5.0
+6K 0.93 0.91 0.92 0.80 0.90 -1.0
+7K 0.90 0.89 0.89 0.84 0.82 6.0
+8K 0.89 0.93 0.88 0.89 0.92 -2.25
+GPT-4o mini 1K 0.85 0.78 0.81 0.80 0.77 4.0
+2K 0.87 0.86 0.87 0.82 0.80 5.5
+3K 0.89 0.88 0.86 0.85 0.80 7.0
+4K 0.79 0.80 0.86 0.83 0.82 0.0
+5K 0.91 0.86 0.88 0.87 0.82 6.0
+6K 0.85 0.89 0.85 0.90 0.85 2.25
+7K 0.92 0.87 0.84 0.90 0.84 4.25
+8K 0.90 0.90 0.84 0.85 0.82 5.25
+Gemma3-4B-Instruct 1K 0.88 0.89 0.85 0.71 0.77 6.25
+2K 0.81 0.91 0.89 0.84 0.83 3.25
+3K 0.80 0.88 0.87 0.92 0.77 9.75
+4K 0.85 0.87 0.92 0.88 0.77 11.0
+5K 0.90 0.83 0.83 0.92 0.82 5.0
+6K 0.93 0.92 0.89 0.93 0.80 11.75
+7K 0.91 0.91 0.90 0.83 0.81 7.75
+8K 0.89 0.88 0.88 0.91 0.84 5.0
+Gemma3-12B-Instruct 1K 0.83 0.84 0.81 0.77 0.78 3.25
+2K 0.83 0.74 0.79 0.85 0.72 8.25
+3K 0.85 0.82 0.77 0.85 0.76 6.25
+4K 0.91 0.87 0.93 0.85 0.84 5.0
+5K 0.85 0.87 0.93 0.83 0.81 6.0
+6K 0.94 0.89 0.93 0.90 0.87 4.5
+7K 0.91 0.85 0.86 0.82 0.77 9.0
+8K 0.90 0.85 0.80 0.85 0.78 7.0
+Table 11: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.78 0.82 0.75 0.75 0.75 2.5
+2K 0.82 0.79 0.75 0.76 0.73 5.0
+3K 0.77 0.81 0.80 0.81 0.71 8.75
+4K 0.79 0.80 0.73 0.77 0.74 3.25
+5K 0.80 0.83 0.83 0.84 0.79 3.5
+6K 0.80 0.82 0.77 0.83 0.72 8.5
+7K 0.84 0.87 0.84 0.79 0.77 6.5
+8K 0.84 0.85 0.82 0.82 0.77 6.25
+Llama3.1-70B-Instruct 1K 0.84 0.81 0.86 0.77 0.75 7.0
+2K 0.79 0.83 0.89 0.84 0.74 9.75
+3K 0.80 0.85 0.83 0.83 0.79 3.75
+4K 0.83 0.81 0.85 0.82 0.74 8.75
+5K 0.84 0.83 0.80 0.77 0.80 1.0
+6K 0.83 0.81 0.80 0.79 0.79 1.75
+7K 0.81 0.75 0.76 0.84 0.77 2.0
+8K 0.87 0.86 0.84 0.85 0.79 6.5
+Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
+2K 0.83 0.89 0.84 0.83 0.82 2.75
+3K 0.83 0.84 0.86 0.79 0.84 -1.0
+4K 0.87 0.83 0.79 0.83 0.84 -1.0
+5K 0.87 0.89 0.89 0.80 0.85 1.25
+6K 0.89 0.81 0.78 0.78 0.82 -0.5
+7K 0.86 0.78 0.79 0.83 0.84 -2.5
+8K 0.86 0.85 0.85 0.85 0.85 0.25
+GPT-4o mini 1K 0.85 0.81 0.81 0.74 0.80 0.25
+2K 0.81 0.80 0.80 0.77 0.75 4.5
+3K 0.83 0.81 0.82 0.74 0.73 7.0
+4K 0.83 0.82 0.79 0.78 0.78 2.5
+5K 0.84 0.82 0.84 0.81 0.75 7.75
+6K 0.85 0.82 0.81 0.79 0.77 4.75
+7K 0.82 0.79 0.81 0.79 0.76 4.25
+8K 0.83 0.83 0.82 0.83 0.78 4.75
+Gemma3-4B-Instruct 1K 0.77 0.82 0.81 0.84 0.73 8.0
+2K 0.86 0.72 0.80 0.85 0.76 4.75
+3K 0.87 0.82 0.84 0.84 0.79 5.25
+4K 0.86 0.85 0.78 0.86 0.80 3.75
+5K 0.88 0.83 0.83 0.84 0.78 6.5
+6K 0.83 0.85 0.86 0.85 0.81 3.75
+7K 0.82 0.82 0.88 0.86 0.82 2.5
+8K 0.87 0.89 0.84 0.83 0.84 1.75
+Gemma3-12B-Instruct 1K 0.90 0.83 0.80 0.82 0.81 2.75
+2K 0.82 0.83 0.87 0.88 0.80 5.0
+3K 0.87 0.84 0.84 0.84 0.76 8.75
+4K 0.80 0.76 0.83 0.79 0.76 3.5
+5K 0.86 0.79 0.86 0.79 0.80 2.5
+6K 0.85 0.87 0.81 0.81 0.81 2.5
+7K 0.88 0.88 0.89 0.86 0.82 5.75
+8K 0.85 0.84 0.85 0.84 0.77 7.5
+Table 12: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
 Models Context Generated Summary Bins Sensitivity
@@
 8K 0.80 0.76 0.73 0.74 0.63 12.75
-Qwen2.5-7B-Instruct 1K 0.78 0.75 0.77 0.72 0.69 6.5
-2K 0.83 0.82 0.74 0.72 0.74 3.75
-3K 0.82 0.78 0.69 0.81 0.72 5.5
-4K 0.85 0.77 0.78 0.72 0.74 4.0
-5K 0.83 0.82 0.79 0.80 0.74 7.0
-6K 0.79 0.75 0.77 0.75 0.70 6.5
-7K 0.80 0.82 0.77 0.80 0.75 4.75
-8K 0.84 0.81 0.78 0.82 0.77 4.25
+Qwen2.5-7B-Instruct 1K 0.84 0.75 0.77 0.76 0.71 7.0
+2K 0.81 0.82 0.78 0.69 0.71 6.5
+3K 0.78 0.76 0.72 0.73 0.68 6.75
+4K 0.90 0.83 0.76 0.80 0.79 3.25
+5K 0.83 0.86 0.74 0.76 0.79 0.75
+6K 0.87 0.82 0.77 0.77 0.83 -2.25
+7K 0.86 0.82 0.79 0.79 0.75 6.5
+8K 0.85 0.88 0.84 0.80 0.76 8.25
 GPT-4o mini 1K 0.79 0.69 0.76 0.73 0.65 9.25
@@
 8K 0.80 0.80 0.76 0.77 0.68 10.25
-Mistral-7B-Instruct-v0.3 1K 0.83 0.77 0.77 0.75 0.80 -2.0
-2K 0.74 0.80 0.80 0.73 0.67 9.75
-3K 0.76 0.75 0.78 0.78 0.70 6.75
-4K 0.79 0.82 0.75 0.80 0.74 5.0
-5K 0.83 0.83 0.84 0.81 0.78 4.75
-6K 0.82 0.81 0.82 0.84 0.79 3.25
-7K 0.83 0.82 0.88 0.87 0.81 4.0
-8K 0.85 0.84 0.83 0.82 0.77 6.5
-Table 8: Comparison of faithfulness scores and Sensitivity for short summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+Gemma3-12B-Instruct 1K 0.75 0.73 0.76 0.75 0.71 3.75
+2K 0.80 0.65 0.70 0.68 0.65 5.75
+3K 0.78 0.76 0.70 0.75 0.66 8.75
+4K 0.82 0.70 0.76 0.76 0.69 7.0
+5K 0.76 0.80 0.86 0.60 0.72 3.5
+6K 0.81 0.79 0.84 0.78 0.69 11.5
+7K 0.83 0.81 0.78 0.79 0.74 6.25
+8K 0.83 0.81 0.78 0.79 0.74 6.25
+Table 13: Comparison of NLI model based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
 The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
@@
 8K 0.74 0.79 0.74 0.77 0.67 9.0
-Qwen2.5-7B-Instruct 1K 0.83 0.76 0.79 0.73 0.70 7.75
-2K 0.81 0.79 0.78 0.78 0.70 9.0
-3K 0.81 0.78 0.80 0.73 0.70 8.0
-4K 0.77 0.76 0.74 0.76 0.69 6.75
-5K 0.82 0.81 0.77 0.76 0.71 8.0
-6K 0.81 0.78 0.78 0.75 0.67 11.0
-7K 0.81 0.74 0.75 0.76 0.75 1.5
-8K 0.79 0.78 0.79 0.78 0.73 5.5
+Qwen2.5-7B-Instruct 1K 0.83 0.74 0.76 0.75 0.68 9.0
+2K 0.76 0.80 0.79 0.76 0.71 6.75
+3K 0.79 0.79 0.81 0.71 0.75 2.5
+4K 0.77 0.75 0.79 0.80 0.73 4.75
+5K 0.83 0.84 0.79 0.76 0.76 4.5
+6K 0.82 0.74 0.76 0.73 0.70 6.25
+7K 0.80 0.74 0.72 0.78 0.78 -2.0
+8K 0.79 0.79 0.80 0.79 0.75 4.25
 GPT-4o mini 1K 0.77 0.74 0.69 0.69 0.69 3.3
@@
 8K 0.76 0.70 0.72 0.69 0.61 10.8
-Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
-2K 0.75 0.80 0.84 0.82 0.81 -0.8
-3K 0.81 0.80 0.82 0.84 0.83 -1.3
-4K 0.79 0.81 0.83 0.81 0.81 0.0
-5K 0.80 0.84 0.86 0.85 0.79 4.8
-6K 0.82 0.87 0.87 0.87 0.86 -0.3
-7K 0.82 0.84 0.82 0.83 0.85 -2.3
-8K 0.87 0.83 0.87 0.89 0.81 5.5
-Table 9: Comparison of faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+Gemma3-12B-Instruct 1K 0.81 0.79 0.72 0.72 0.73 3.0
+2K 0.77 0.75 0.79 0.78 0.74 3.25
+3K 0.75 0.77 0.71 0.73 0.68 6.0
+4K 0.72 0.67 0.70 0.74 0.65 5.75
+5K 0.77 0.78 0.80 0.75 0.71 6.5
+6K 0.77 0.78 0.77 0.74 0.70 6.5
+7K 0.82 0.81 0.82 0.79 0.71 10.0
+8K 0.74 0.74 0.77 0.75 0.63 12.0
+Table 14: Comparison of NLI model based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
+Models Context Generated Summary Bins
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.24 0.29 0.14 0.24 0.33
+2K 0.21 0.21 0.16 0.24 0.37
+3K 0.20 0.30 0.20 0.10 0.57
+4K 0.06 0.17 0.23 0.17 0.26
+5K 0.31 0.20 0.20 0.36 0.33
+6K 0.21 0.25 0.29 0.14 0.50
+7K 0.24 0.17 0.14 0.31 0.66
+8K 0.28 0.24 0.31 0.24 0.69
+Qwen2.5-7B-Instruct 1K 0.07 0.00 0.13 0.00 0.33
+2K 0.13 0.20 0.13 0.00 0.13
+3K 0.27 0.20 0.27 0.20 0.00
+4K 0.27 0.00 0.20 0.13 0.20
+5K 0.13 0.07 0.33 0.20 0.27
+6K 0.07 0.07 0.27 0.13 0.20
+7K 0.31 0.00 0.08 0.00 0.00
+8K 0.31 0.23 0.23 0.23 0.23
+GPT-4o mini 1K 0.05 0.00 0.00 0.00 0.05
+2K 0.00 0.00 0.10 0.00 0.10
+3K 0.05 0.00 0.00 0.00 0.10
+4K 0.00 0.00 0.05 0.05 0.15
+5K 0.15 0.05 0.05 0.05 0.10
+6K 0.10 0.10 0.00 0.05 0.15
+7K 0.10 0.00 0.00 0.05 0.10
+8K 0.15 0.15 0.10 0.20 0.20
+Gemma3-12B-Instruct 1K 0.12 0.00 0.12 0.00 0.12
+2K 0.25 0.25 0.00 0.12 0.12
+3K 0.00 0.00 0.00 0.12 0.00
+4K 0.00 0.00 0.00 0.11 0.00
+5K 0.12 0.00 0.00 0.00 0.12
+6K 0.00 0.00 0.00 0.00 0.14
+7K 0.00 0.00 0.00 0.00 0.33
+8K 0.00 0.00 0.00 0.00 0.17
+Table 15: Comparison of FineSurE error rates for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest error rate in each bin is marked in bold, while the lowest is underlined.
+Models Context Generated Summary Bins
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.53 0.27 0.60 0.20 0.80
+2K 0.80 0.53 0.67 0.80 1.47
+3K 1.75 1.12 1.12 1.00 2.25
+4K 0.89 1.22 1.44 0.44 2.44
+5K 0.60 1.07 1.13 0.87 2.07
+6K 0.83 0.67 0.67 0.67 1.50
+7K 0.90 0.60 1.00 0.90 1.90
+8K 0.67 0.92 0.75 0.83 1.17
+Qwen2.5-7B-Instruct 1K 0.23 0.31 0.23 0.15 0.15
+2K 0.31 0.00 0.54 0.38 0.23
+3K 0.25 0.25 0.42 0.33 0.17
+4K 0.50 0.20 0.20 0.50 0.70
+5K 0.20 0.30 0.00 0.40 0.30
+6K 0.25 0.92 0.67 0.67 0.33
+7K 0.64 0.82 0.82 0.64 0.27
+8K 0.10 0.10 0.30 0.30 1.2
+GPT-4o mini 1K 0.20 0.25 0.10 0.20 0.15
+2K 0.10 0.20 0.30 0.15 0.40
+3K 0.43 0.36 0.57 0.21 0.64
+4K 0.67 0.28 0.39 0.17 0.72
+5K 0.31 0.50 0.25 0.12 0.31
+6K 0.39 0.61 0.56 0.50 0.72
+7K 0.50 0.71 0.36 0.36 0.50
+8K 0.67 0.75 0.92 0.92 0.75
+Gemma3-12B-Instruct 1K 0.00 0.00 0.00 0.00 0.20
+2K 0.00 0.00 0.00 0.00 0.12
+3K 0.30 0.30 0.00 0.30 0.30
+4K 0.40 0.60 0.80 1.00 0.40
+5K 0.25 0.00 0.25 0.25 0.75
+6K 0.12 0.12 0.25 0.38 0.12
+7K 0.00 0.00 0.00 0.00 0.14
+8K 0.00 0.17 0.00 0.33 0.00
+Table 16: Comparison of FineSurE error rates for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest error rate in each bin is marked in bold, while the lowest is underlined.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.66 0.63 0.63 0.62 0.57 6.5
+2K 0.70 0.62 0.64 0.60 0.57 7.0
+3K 0.70 0.64 0.63 0.63 0.61 4.0
+4K 0.60 0.66 0.63 0.61 0.58 4.5
+5K 0.70 0.65 0.66 0.64 0.61 5.25
+6K 0.70 0.67 0.69 0.66 0.62 6.0
+7K 0.74 0.67 0.64 0.65 0.60 7.5
+8K 0.68 0.65 0.60 0.64 0.55 9.25
+Qwen2.5-7B-Instruct 1K 0.70 0.58 0.65 0.61 0.58 5.5
+2K 0.81 0.61 0.64 0.57 0.52 13.75
+3K 0.70 0.62 0.57 0.60 0.54 8.25
+4K 0.73 0.68 0.63 0.59 0.57 8.75
+5K 0.71 0.65 0.66 0.57 0.57 7.75
+6K 0.70 0.60 0.62 0.62 0.58 5.5
+7K 0.72 0.57 0.61 0.61 0.54 8.75
+8K 0.66 0.60 0.64 0.61 0.59 3.75
+GPT-4o mini 1K 0.68 0.60 0.55 0.51 0.51 7.5
+2K 0.73 0.58 0.58 0.57 0.50 11.5
+3K 0.72 0.59 0.58 0.57 0.53 8.5
+4K 0.68 0.60 0.55 0.56 0.52 7.75
+5K 0.68 0.61 0.60 0.56 0.52 9.25
+6K 0.67 0.61 0.60 0.57 0.52 9.25
+7K 0.68 0.62 0.58 0.57 0.53 8.25
+8K 0.69 0.62 0.56 0.56 0.49 11.75
+Gemma3-12B-Instruct 1K 0.64 0.55 0.56 0.51 0.62 -5.5
+2K 0.65 0.56 0.52 0.53 0.54 2.5
+3K 0.65 0.63 0.55 0.55 0.53 6.5
+4K 0.63 0.59 0.57 0.56 0.52 6.75
+5K 0.61 0.57 0.62 0.57 0.53 6.25
+6K 0.69 0.63 0.60 0.56 0.54 8.0
+7K 0.68 0.61 0.63 0.67 0.56 8.75
+8K 0.68 0.61 0.63 0.67 0.56 8.75
+Table 17: Comparison of ROUGE-L based faithfulness scores and Sensitivity for standard summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
+Models Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Llama3.1-8B-Instruct 1K 0.65 0.62 0.60 0.61 0.58 4.0
+2K 0.65 0.66 0.64 0.62 0.57 7.25
+3K 0.65 0.66 0.65 0.64 0.56 9.0
+4K 0.62 0.61 0.60 0.64 0.60 1.75
+5K 0.65 0.66 0.65 0.63 0.58 6.75
+6K 0.65 0.71 0.71 0.69 0.61 8.0
+7K 0.68 0.68 0.64 0.64 0.60 6.0
+8K 0.62 0.67 0.67 0.67 0.62 3.75
+Qwen2.5-7B-Instruct 1K 0.69 0.57 0.58 0.62 0.52 9.5
+2K 0.62 0.62 0.60 0.60 0.54 7.0
+3K 0.66 0.60 0.60 0.62 0.56 6.0
+4K 0.63 0.61 0.61 0.66 0.57 5.75
+5K 0.64 0.65 0.63 0.63 0.60 3.75
+6K 0.65 0.63 0.64 0.61 0.57 6.25
+7K 0.65 0.66 0.62 0.60 0.57 6.25
+8K 0.64 0.66 0.63 0.61 0.57 6.5
+GPT-4o mini 1K 0.69 0.57 0.55 0.58 0.52 7.75
+2K 0.61 0.55 0.54 0.53 0.50 5.75
+3K 0.59 0.55 0.55 0.50 0.47 7.75
+4K 0.62 0.58 0.56 0.55 0.53 4.75
+5K 0.59 0.55 0.55 0.52 0.46 9.25
+6K 0.59 0.56 0.55 0.53 0.49 6.75
+7K 0.58 0.56 0.53 0.51 0.49 5.5
+8K 0.58 0.54 0.53 0.50 0.47 6.75
+Gemma3-12B-Instruct 1K 0.66 0.57 0.60 0.61 0.47 14.0
+2K 0.63 0.60 0.60 0.61 0.57 4.0
+3K 0.59 0.59 0.54 0.57 0.56 1.25
+4K 0.61 0.56 0.58 0.58 0.58 0.25
+5K 0.62 0.60 0.61 0.63 0.54 7.5
+6K 0.60 0.63 0.61 0.59 0.57 3.75
+7K 0.65 0.65 0.60 0.61 0.57 5.75
+8K 0.60 0.64 0.60 0.60 0.56 5.0
+Table 18: Comparison of ROUGE-L based faithfulness scores and Sensitivity for long summaries generated by different models across output bins on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
+Datasets Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Wikipedia 1K 0.81 0.79 0.85 0.82 0.79 2.75
+2K 0.84 0.79 0.81 0.81 0.77 4.25
+3K 0.87 0.85 0.84 0.83 0.79 5.75
+4K 0.91 0.90 0.88 0.86 0.88 0.75
+5K 0.85 0.85 0.86 0.83 0.83 1.75
+6K 0.89 0.85 0.85 0.83 0.81 4.5
+7K 0.91 0.81 0.85 0.87 0.81 5.0
+8K 0.87 0.84 0.82 0.82 0.79 4.75
+Arxiv 1K 0.71 0.79 0.72 0.80 0.69 6.5
+2K 0.82 0.71 0.76 0.69 0.75 -0.5
+3K 0.74 0.87 0.78 0.82 0.77 3.25
+4K 0.87 0.86 0.84 0.85 0.79 6.5
+5K 0.84 0.85 0.79 0.82 0.81 1.5
+6K 0.82 0.83 0.77 0.85 0.90 -8.25
+7K 0.86 0.82 0.89 0.84 0.79 6.25
+8K 0.80 0.81 0.75 0.78 0.81 -2.5
+Pubmed 1K 0.88 0.87 0.88 0.79 0.86 -0.5
+2K 0.89 0.85 0.81 0.87 0.85 0.5
+3K 0.91 0.89 0.86 0.91 0.90 -0.75
+4K 0.93 0.90 0.92 0.87 0.92 -1.5
+5K 0.91 0.84 0.89 0.91 0.85 3.75
+6K 0.91 0.88 0.94 0.91 0.91 0.0
+7K 0.91 0.89 0.92 0.91 0.92 -1.25
+8K 0.91 0.92 0.89 0.87 0.90 -0.25
+Govreport 1K 0.87 0.82 0.86 0.83 0.81 3.5
+2K 0.94 0.88 0.90 0.93 0.93 -1.75
+3K 0.91 0.92 0.90 0.91 0.91 0.0
+4K 0.94 0.83 0.88 0.85 0.84 3.5
+5K 0.94 0.93 0.94 0.93 0.93 0.5
+6K 0.93 0.91 0.92 0.94 0.87 5.5
+7K 0.91 0.89 0.94 0.97 0.89 3.75
+8K 0.90 0.93 0.93 0.90 0.93 -1.5
+Table 19: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
+The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
+Datasets Context Generated Summary Bins Sensitivity
+Length 1 2 3 4 5
+Wikipedia 1K 0.78 0.82 0.75 0.75 0.75 2.5
+2K 0.82 0.79 0.75 0.76 0.73 5.0
+3K 0.77 0.81 0.80 0.81 0.71 8.75
+4K 0.79 0.80 0.73 0.77 0.74 3.25
+5K 0.80 0.83 0.83 0.84 0.79 3.5
+6K 0.80 0.82 0.77 0.83 0.72 8.5
+7K 0.84 0.87 0.84 0.79 0.77 6.5
+8K 0.84 0.85 0.82 0.82 0.77 6.25
+Arxiv 1K 0.73 0.81 0.81 0.73 0.72 5.0
+2K 0.80 0.75 0.78 0.72 0.75 1.25
+3K 0.78 0.77 0.76 0.75 0.72 4.5
+4K 0.74 0.80 0.70 0.78 0.74 1.5
+5K 0.81 0.80 0.84 0.83 0.77 5.0
+6K 0.70 0.77 0.72 0.72 0.74 -1.25
+7K 0.86 0.81 0.76 0.74 0.71 8.25
+8K 0.79 0.74 0.81 0.77 0.68 9.75
+Pubmed 1K 0.89 0.87 0.93 0.87 0.89 0.0
+2K 0.84 0.87 0.81 0.88 0.81 4.0
+3K 0.86 0.89 0.90 0.90 0.87 1.75
+4K 0.83 0.89 0.84 0.86 0.79 6.5
+5K 0.87 0.87 0.86 0.85 0.85 1.25
+6K 0.87 0.90 0.86 0.81 0.80 6.0
+7K 0.91 0.89 0.87 0.82 0.82 5.25
+8K 0.90 0.82 0.86 0.81 0.71 13.75
+Govreport 1K 0.88 0.86 0.86 0.89 0.85 2.25
+2K 0.89 0.90 0.88 0.91 0.91 -1.5
+3K 0.87 0.92 0.91 0.91 0.88 2.25
+4K 0.90 0.89 0.94 0.90 0.82 8.75
+5K 0.94 0.91 0.90 0.91 0.89 2.5
+6K 0.91 0.91 0.90 0.92 0.92 -1.0
+7K 0.93 0.94 0.93 0.93 0.91 2.25
+8K 0.92 0.93 0.95 0.92 0.85 8.0
+Table 20: Comparison of MiniCheck-Flan-T5-Large based faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
 The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
@@
 8K 0.84 0.82 0.87 0.85 0.81 3.5
-Table 10: Comparison of faithfulness scores and Sensitivity for short summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
+Table 21: Comparison of NLI model based faithfulness scores and Sensitivity for standard summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
 The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.
@@
 8K 0.84 0.87 0.89 0.84 0.80 6.0
-Table 11: Comparison of faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
+Table 22: Comparison of NLI model based faithfulness scores and Sensitivity for long summaries generated by the Llama3.1-8B-Instruct model across output bins in multiple domains, with context lengths ranging from 1K to 8K.
 The highest faithfulness score in each bin is marked in bold, while the lowest is underlined.

### v3

@@
 While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs.
-In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study.
+In this work, we investigate where hallucinations frequently occur in LLM-based long response generation, using long document summarization as a key case study.
 Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response.
-To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences.
+To understand this positional bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences.
 Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
@@
 (2024).
-Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
+Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words using GPT-4o mini.
 As LLMs continue to evolve, a paradigm shift is emerging from short-form to long-form generation, enabled by the ability to process extended contexts Wu et al.
@@
 2 Generating Summaries
-This section describes the experimental setup used to generate the summaries analyzed in our study.
-We prompt LLMs to generate summaries from original documents, varying both the context length and the output length.
-For all experiments, we generate summaries using greedy decoding.
-The prompt is provided in Appendix A.
+This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents. The prompt is provided in Appendix G.
 2.1 Input Context Lengths
-To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 8K tokens.
-For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
+We evaluate model performance with input length of over 4K tokens. For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
 2.2 Output Length Categories
@@
 By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases.
-See Appendix A for more details.
+See Appendix G for more details.
 2.3 Overall Faithfulness
 As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al.
-(2024) to ensure consistent quality across different context lengths, prior to our main positional analysis.
+(2024) to ensure consistent quality over various context lengths, prior to our main positional analysis.
 Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al.
-(2024) on the Wikipedia dataset (1K–8K context lengths), evaluated using FineSurE with GPT-4 Achiam et al.
+(2024) on the Wikipedia dataset, evaluated using FineSurE with GPT-4 Achiam et al.
 (2023).
@@
 We next conduct a finer-grained positional analysis in Section 3 for various setups.
+3 Where Do Hallucinations Most Frequently Occur?
+In this section, we empirically investigate where hallucinations most frequently occur in long document summarization.
+We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
 Context Length Faithfulness(%)
-1K 83.3
-2K 86.7
-3K 78.9
 4K 82.9
@@
 8K 86.8
-Table 1: Faithfulness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
-3 Where Do Hallucinations Most Frequently Occur?
-In this section, we empirically investigate where hallucinations most frequently occur in long document summarization.
-We analyze the positional distribution of factual errors by examining faithfulness across varying models, datasets, and output positions.
-Figure 2: Comparison of faithfulness scores for standard summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
+Table 1: Overall faithfulness of long summaries on the Wikipedia dataset, evaluated using FineSurE.
 3.1 Experimental Setup
@@
 (2024a), and GPT-4o mini, to ensure model-agnostic findings.
-For Gemma, we retain the default sliding window attention context length of 1024 tokens, while for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
+For Gemma, we set the sliding window attention context length to 1 token (as it cannot be disabled entirely), whereas for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
+Throughout all experiments, summaries are generated using greedy decoding.
 Datasets
@@
 (2021) datasets for further evaluation.
+For each dataset, we randomly sampled 20 examples for analysis.
 3.2 Evaluation Metric
@@
 (2024) model to evaluate faithfulness.
-MiniCheck is particularly suitable for this granular task, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al.
+MiniCheck is particularly suitable for this fine-grained analysis, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al.
 (2023); Jing et al.
@@
 (2024).
-Let $s_{i}$ be the $i$-th sentence in the generated summary $S$.
-Each sentence in the generated summary is decomposed into a set of atomic facts $A_{i}$ by utilizing LLM, followed by a filtering process to remove unnecessary atomic facts Yang et al.
+Let $s_{i}$ denote the $i$-th sentence in a generated summary $S$.
+Each sentence is first decomposed into a set of atomic facts $A_{i}$ using an LLM, followed by a filtering step to remove unnecessary atomic facts Yang et al.
 (2024b), which makes:
 $A_{i}=\{a_{i1},a_{i2},...,a_{iN_{i}}\}$ (1)
-Using a MiniCheck model, which is denoted as $\mathcal{M}$, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner.
-Each atomic fact is assigned the highest score among its comparisons:
+We then compare each filtered atomic fact with the source document $D={d_{1},...,d_{M}}$ using the MiniCheck model $\mathcal{M}$ in a pairwise fashion.
+For each atomic fact, we take the maximum score obtained across all source sentences:
 $\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\mathcal{M}(d_{m},a_{ij})$ (2)
-Finally, we use the average of the scores for a set of filtered atomic facts in a sentence as its faithfulness score:
+Finally, the sentence-level faithfulness score is computed as the average score of all filtered atomic facts within that sentence:
 $\textrm{Faithfulness}(s_{i})=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\textrm{score}(a_{ij})$ (3)
-Additional experimental details and an example of the evaluation procedure are provided in Appendix B.
-Furthermore, we verify the quality of atomic facts generated by the LLM via human evaluation in Table 2 for atomic facts extracted from long summaries generated by Llama.
-Context Length Validation(%)
-1K 98.55
-2K 99.08
-3K 97.62
-4K 95.92
-5K 99.01
-6K 96.18
-7K 99.12
-8K 99.64
-Table 2: Human evaluation of the quality (completeness) of atomic facts.
-For each context length, we randomly sampled 10 summaries and conducted the evaluation manually.
-Each atomic fact was annotated as either true or false depending on whether it is present in the summary.
+Each sentence is assigned a faithfulness score in the range $[0,1]$, where 1 indicates that all atomic facts in the sentence are fully supported by the source document, and 0 indicates a complete lack of support.
 Sensitivity
 In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs.
-Specifically, following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins or segments, and the average faithfulness score of the atomic facts within each bin is computed.
-We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins.
-A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise.
-The larger the sensitivity, the more pronounced the hallucination at the end of the output.
+Following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed.
+The bins are defined using the cut function in the pandas package, and the scores of the atomic facts are assigned to bins using open-interval boundaries, meaning that boundary values are not included.
+Let $f_{k}$ denote the average faithfulness score of the $k$-th bin ($k=1,\dots,5$).
+We define sensitivity as:
+$\text{Sensitivity}=100\times\left(\frac{1}{4}\sum_{k=1}^{4}f_{k}-f_{5}\right).$
+A positive value indicates that the faithfulness in the final bin is lower than in the earlier bins (i.e., a tendency to hallucinate at the end), whereas a negative value suggests the opposite trend.
+Larger sensitivity values correspond to stronger hallucination effects near the end of the generated output.
+Additional experimental details and an example of the evaluation procedure are provided in Appendix C.
+Furthermore, the human evaluation of the quality and completeness of the generated atomic facts and experimental details is presented in Appendix B.
+3.3 Hallucinate at the Last
+The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across summary lengths.
+The decline becomes most pronounced when the summaries are long.
+We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset in Figure 2.
+As shown in the figure, both standard and long summaries produced by the three models, excluding Qwen, consistently exhibit a decline in faithfulness toward the later portions of the summary, with the lowest scores typically appearing at the end.
+This Hallucinate at the Last phenomenon becomes even more pronounced in long summaries: as the summary length increases, the faithfulness near the end continues to decrease, eventually dropping below 0.75 for these models.
+In contrast, Qwen displays a different pattern.
+Its faithfulness tends to drop most sharply in the middle of the summary, after which the scores recover toward the end.
+Figure 2: Comparison of faithfulness scores for summaries generated by different models on the Wikipedia dataset.
+Models Generated Summary Bins Sensitivity
+1 2 3 4 5
+Llama3.1-8B-Instruct 0.81 0.82 0.79 0.80 0.75 5.50
+Llama3.1-70B-Instruct 0.83 0.82 0.83 0.81 0.77 5.25
+Gemma3-12B-Instruct 0.85 0.83 0.84 0.83 0.79 4.75
+GPT-4o mini 0.83 0.81 0.81 0.78 0.76 4.75
+Qwen2.5-7B-Instruct 0.86 0.83 0.83 0.82 0.84 -0.50
+Table 2: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset.
 Models Context Generated Summary Bins Sensitivity
 Length 1 2 3 4 5
-Llama3.1-8B-Instruct Overall Avg. 0.81 0.82 0.79 0.80 0.75 5.5
-12K 0.80 0.79 0.66 0.63 0.69 9.5
+Standard Summary
+Llama3.1-8B-Instruct 12K 0.91 0.87 0.85 0.76 0.67 17.65
+16K 0.84 0.83 0.76 0.71 0.69 3.00
+Long Summary
+Llama3.1-8B-Instruct 12K 0.80 0.79 0.66 0.63 0.69 9.50
 16K 0.86 0.83 0.85 0.83 0.79 5.25
-Llama3.1-70B-Instruct Overall Avg. 0.83 0.82 0.83 0.81 0.77 5.25
-Gemma3-4B-Instruct Overall Avg. 0.85 0.82 0.83 0.85 0.79 4.75
-Gemma3-12B-Instruct Overall Avg. 0.85 0.83 0.84 0.83 0.79 4.75
-GPT-4o mini Overall Avg. 0.83 0.81 0.81 0.78 0.76 4.75
-Qwen2.5-7B-Instruct Overall Avg. 0.86 0.83 0.83 0.82 0.84 -0.5
-Table 3: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset.
-Overall Avg.
-denotes the average faithfulness score aggregated across context lengths from 1K to 8K.
-3.3 Hallucinate at the Last
-Model Context Generated Summary Bins Sensitivity
-Length 1 2 3 4 5
-Llama3.1-8B-Instruct 1K 0.92 0.90 0.91 0.87 0.68 22.00
-2K 0.82 0.74 0.83 0.72 0.71 6.75
-3K 0.94 0.94 0.81 0.86 0.68 20.75
-4K 0.92 0.92 0.88 0.87 0.82 7.75
-5K 0.91 0.88 0.89 0.78 0.70 16.50
-6K 0.86 0.84 0.79 0.89 0.73 11.50
-7K 0.97 0.89 0.82 0.94 0.66 24.50
-8K 0.93 0.90 0.86 0.79 0.65 22.00
+Table 3: Faithfulness scores and sensitivity for standard and long summaries generated by Llama on the Wikipedia dataset, with 12K and 16K context length.
+Model Generated Summary Bins Sensitivity
+1 2 3 4 5
+Llama3.1-8B-Instruct 0.90 0.87 0.84 0.84 0.70 16.25
 Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset.
-Each atomic fact was annotated as either true or false depending on whether it is present in the document.
-The results were manually evaluated and aggregated.
-The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across various contexts and summary lengths.
-The decline becomes most pronounced when the context length reaches 8K and the summaries are long.
-We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively.
-As shown in Figure 2, standard summaries generated by the three models, excluding Qwen, consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end.
-Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length.
-Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
-Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
-However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries except for Qwen, as shown in Figure 3.
-For these models, as the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below a faithfulness score of 0.70 when 45 sentences are generated from an 8K context.
-Qwen, in contrast, often exhibits the largest drop in faithfulness in the middle of the summary, after which the scores recover toward the end.
-Tables 3 report the faithfulness and sensitivity of various models.
-We aggregate the faithfulness scores across context lengths from 1K to 8K and report the overall average.
-All models except Qwen show markedly low faithfulness scores in the final bin, reflecting high sensitivity.
-Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size or context length.
-Qwen, however, is the only model that records sensitivity values below zero.
-Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity, and we provide a detailed analysis of these contrasting patterns in Section 4.
-Supplementary results across context lengths and methods are provided in Appendix C.
-The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, with a marked decline once the context length exceeds 7K.
+Figure 3: Comparison of faithfulness scores for summaries generated by the Llama3.1-8B-Instruct model in multiple domains.
+Table 2 report the faithfulness and sensitivity values for the different models.
+All models except Qwen show markedly low faithfulness scores in the final bin, indicating high sensitivity.
+Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size.
+In contrast, Qwen is the only model that records negative sensitivity values.
+Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity.
+We provide a detailed analysis of these contrasting patterns in Section 4.
+Table 3 presents the faithfulness and sensitivity results for both the standard and long summaries generated by the LLaMA model with context lengths of 12K and 16K tokens.
+Interestingly, we find that positional discrepancies in the generated summaries are not significantly affected by the input context length.
+As shown in the table, even with longer contexts of approximately 12K and 16K tokens, sensitivity remains high, consistently demonstrating the Hallucinate at the Last phenomenon.
+Supplementary results across different context lengths and methods are presented in Appendix D.
+The latter part of the summary consistently demonstrates the lowest level of faithfulness across all datasets, with a marked decline on the long summaries.
 We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 3.
 As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains.
-Faithfulness scores tend to decrease toward the final segments as the output length increases.
-In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline, with the deterioration becoming especially pronounced once the context length exceeds 7K.
-Additional sensitivity results across diverse datasets are provided in Appendix D.
-Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
+Faithfulness scores gradually decrease toward the final segments as the output length increases, and this decline becomes most severe in the concluding portions of longer summaries.
+Additional sensitivity results across diverse datasets are provided in Appendix E.
+The Hallucinate at the Last phenomenon persists across different decoding strategies.
+To demonstrate that the Hallucinate at the Last phenomenon is not merely a consequence of error propagation, we first conduct experiments using various decoding strategies.
+Following  Shi et al.
+(2024a), we adopt the following decoding strategies to generate summaries.
+  • Temperature & Top-p Sampling samples tokens from the next-token probability distribution using the smallest set of tokens whose cumulative probability mass exceeds a given threshold p.
+We conduct experiments with temperature values $\tau\in\{0.5,0.7\}$ and with top-p thresholds of 0.7 and 0.9.
+  • Top-k Sampling samples tokens from the k most probable candidates. We experiment with k values of 20 and 100.
+  • $\eta$-Sampling discards tokens whose probabilities fall below an entropy dependent threshold. The hyperparameter $\eta$ is explored with 6e-4 and 4e-3.
+We present the positional faithfulness discrepancies in long summaries generated using different decoding strategies in Figure 4.
+For the Llama and Gemma models, the latter portions of the summaries exhibit the lowest levels of faithfulness across all decoding strategies.
+This effect is particularly pronounced under $\eta$-sampling for Llama, where the faithfulness score remains above 0.8 in the earlier segments but drops below 0.7 in the final bin.
+In contrast, the Qwen model shows a different trend.
+With the exception of two decoding strategies, the faithfulness of the generated summaries does not decline toward the end but instead increases.
+Notably, the faithfulness scores for Qwen are lowest in the middle portions of the summaries rather than at the end.
+Motivated by these observations, we further investigate this phenomenon across a broader set of models.
+Figure 4: Comparison of faithfulness scores for long summaries generated using different decoding strategies on the Wikipedia dataset.
+Human Evaluation
+Each atomic fact was labeled as true or false based on whether it was supported by the source document.
+True labels were scored as 1 and false as 0, and these scores were averaged within the same five positional bins as in the automatic evaluation.
+The resulting human-judged faithfulness scores are reported in Table 4.
+Furthermore, to assess the reliability of the human evaluation, we measured inter-annotator agreement among annotators who evaluated the same subset.
+In total, 543 atomic facts were evaluated, and the annotators agreed on 515 of them, corresponding to a 94.8% raw agreement.
 4 Why Do Hallucinations Frequently Occur at the Last?
@@
 Experimental Setup
-We use human-written reference summaries from the CNNDM dataset, selecting examples with long summaries and input contexts of around 2K tokens.
-We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
+We use human-written reference summaries from the CNNDM dataset, selecting examples with long reference summaries. We compare these references to summaries generated by the Llama3.1-8B-Instruct model.
 Results & Analysis
 Figure 5 compares positional faithfulness between human-written and model-generated CNNDM summaries.
-We observe that the faithfulness of the generated summaries declines steadily, falling below 0.6 in the final segment, whereas reference summaries dip around the middle but recover.
+We observe that the faithfulness of the generated summaries declines steadily, falling below 0.65 in the final segment, whereas reference summaries dip around the middle but recover.
 This suggests that Hallucination at the Last is not solely due to summarization’s inherent structure.
@@
 Experimental Setup
-Unlike previous studies that focus on attention to the first generated token Hsieh et al.
-(2024), we compute attention weights at the sentence level.
-Moreover, we segment the full sequence into chunks of 100 tokens.
-Our analysis specifically focuses on three positions within the generated output: the first sentence, the middle sentence, and the final sentence.
-Details of the attention weight calculation are provided in Appendix E.
+Following Hsieh et al.
+(2024), we compute attention weights by averaging the decoder self-attention matrices across all layers and heads.
+Unlike the prior study, however, we compute sentence-level attention by aligning token spans with sentence boundaries and averaging the token-level attention weights within each sentence.
+We segment the sequence into 100-token chunks and focus our analysis on three positions: the first, middle, and final sentences.
+Further details are provided in Appendix A.
 Results & Analysis
 Figure 6 presents a visualization of average attention weights across sentences.
-As shown in the results, the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 3, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences.
-In contrast, the Qwen model assigns similar levels of attention to all three sentence positions.
+As shown in Figure 6(a), the Llama model, which exhibited the Hallucinate at the Last pattern in Figure 2, assigns nearly three times more attention to the final sentence of the generated summary compared to the first and middle sentences.
+In contrast, the Qwen model (Figure 6(b)) assigns similar levels of attention to all three sentence positions.
 These findings suggest that increased attention to previously generated text correlates with a higher likelihood of hallucination.
@@
 (2019), which maintains a balanced focus across the context and mitigates the downstream attention concentration that triggers hallucinations.
-Table 5 compares the Qwen model with and without sliding window attention.
-In this setting, the sliding window context length is fixed at 16000 tokens.
-The results clearly show that sliding window attention omission increases sensitivity across all context and summary lengths, up to 9.75.
-They demonstrate that employing sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
+Table 5 compares Qwen with and without sliding window attention, where the sliding-window context length is fixed at 16,000 tokens.
+The results clearly indicate that removing the sliding window attention mechanism increases the sensitivity to 3.75 for standard summaries and 3.25 for long summaries.
+Furthermore, as shown in Figure 6(c), the attention level becomes higher than that observed with sliding window attention.
+While autoregressive decoders generally allocate more attention to recent tokens, this trend is not uniform across models.
+Taken together, these results demonstrate that sliding window attention not only enhances decoding efficiency but also reduces attention concentration on later tokens as the generated sequence length increases.
 (a) Average attention weights of summaries generated by Llama.
-(b) Average attention weights of summaries generated by Qwen.
+(b) Average attention weights of summaries generated by Qwen with sliding window attention.
+(c) Average attention weights of summaries generated by Qwen without sliding window attention.
 Figure 6: Average attention weights of two models exhibiting contrasting trends in Figure 2. The dashed line separates the input context from the generated output.
+Models Generated Summary Bins Sensitivity
+1 2 3 4 5
+Standard Summary
+Qwen2.5-7B-Instruct 0.89 0.88 0.85 0.81 0.87 -1.25
+w/o sliding window attention 0.90 0.88 0.85 0.83 0.84 2.5(+3.75)
+Long Summary
+Qwen2.5-7B-Instruct 0.86 0.83 0.83 0.82 0.84 -0.50
+w/o sliding window attention 0.86 0.84 0.83 0.82 0.81 2.75(+3.25)
+Table 5: Faithfulness scores and sensitivity for standard and long summaries generated by Qwen2.5-7B-Instruct on the Wikipedia dataset, with and without sliding window attention.
+Note that higher sensitivity indicates stronger hallucination effects toward the end of the output.
 5 How to Mitigate the Hallucinate at the Last?
@@
 To this end, we apply four methods that may mitigate the phenomenon.
-Models Context Generated Summary Bins Sensitivity
-Length 1 2 3 4 5
-Standard Summary
-Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
-2K 0.90 0.89 0.83 0.78 0.84 1.0
-3K 0.87 0.84 0.79 0.83 0.82 1.25
-4K 0.91 0.88 0.90 0.79 0.89 -2.0
-5K 0.86 0.89 0.85 0.76 0.89 -5.0
-6K 0.93 0.91 0.92 0.80 0.90 -1.0
-7K 0.90 0.89 0.89 0.84 0.82 6.0
-8K 0.89 0.93 0.88 0.89 0.92 -2.25
-Without 1K 0.88 0.79 0.74 0.76 0.84 -4.75(+1.5)
-2K 0.90 0.89 0.83 0.78 0.84 1.0(+0.0)
-Sliding 3K 0.88 0.86 0.82 0.85 0.83 2.25(+1.0)
-4K 0.91 0.88 0.87 0.83 0.83 4.25(+6.25)
-Winodw 5K 0.87 0.87 0.81 0.82 0.80 4.25(+9.25)
-6K 0.93 0.90 0.91 0.85 0.86 3.75(+4.75)
-Attention 7K 0.91 0.91 0.89 0.84 0.79 9.75(+3.75)
-8K 0.88 0.93 0.89 0.89 0.92 -2.25(+0.0)
-Long Summary
-Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
-2K 0.83 0.89 0.84 0.83 0.82 2.75
-3K 0.83 0.84 0.86 0.79 0.84 -1.0
-4K 0.87 0.83 0.79 0.83 0.84 -1.0
-5K 0.87 0.89 0.89 0.80 0.85 1.25
-6K 0.89 0.81 0.78 0.78 0.82 -0.5
-7K 0.86 0.78 0.79 0.83 0.84 -2.5
-8K 0.86 0.85 0.85 0.85 0.85 0.25
-Without 1K 0.89 0.77 0.83 0.84 0.80 3.25(+4.25)
-2K 0.83 0.89 0.83 0.80 0.78 5.75(+3.0)
-Sliding 3K 0.83 0.84 0.86 0.79 0.84 -1.0(+0.0)
-4K 0.87 0.83 0.79 0.83 0.84 -1.0(+0.0)
-Winodw 5K 0.86 0.90 0.83 0.79 0.80 4.5(+3.25)
-6K 0.89 0.82 0.84 0.83 0.78 6.5(+7.0)
-Attention 7K 0.86 0.80 0.78 0.83 0.82 -0.25(+2.25)
-8K 0.88 0.85 0.85 0.86 0.82 4.0(+3.75)
-Table 5: Faithfulness scores and sensitivity for standard and long summaries generated by Qwen2.5-7B-Instruct on the Wikipedia dataset, with and without sliding window attention.
-Note that higher sensitivity indicates stronger hallucination effects toward the end of the output.
 Experimental Setup
-In this experiment, we generate Wikipedia summaries with a 7K context using four methods, taking Llama3.1-8B-Instruct as the baseline.
+In this experiment, we generate Wikipedia summaries using four methods, taking Llama3.1-8B-Instruct as the baseline.
 For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
@@
   • LongWriter-Llama3.1-8B Bai et al. (2025) is a model fine-tuned on a long-output dataset and further enhanced using DPO Rafailov et al. (2023).
-  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024).
+  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024b).
 Methods Generated Summary Bins Sensitivity
@@
 Llama3.1-8B 0.74 0.75 0.75 0.73 0.64 10.3
-+ BooookScore 0.73 0.77 0.78 0.80 0.75 2.0
++ BooookScore 0.73 0.77 0.78 0.80 0.77 0.0
 + MInference 0.77 0.78 0.78 0.71 0.69 7.0
@@
 Table 6: Faithfulness scores and sensitivity for long summaries generated by different mitigation methods on the Wikipedia dataset.
+Figure 7: Faithfulness scores for long summaries generated by different mitigation methods on the Wikipedia dataset across increasing output lengths.
 Results & Analysis
-We report the faithfulness scores across output bins and the sensitivity results in Table 6, and the faithfulness scores across increasing output lengths for each method in Figure 7.
+We report the faithfulness scores across output bins and the corresponding sensitivity values in Table 6, and present the bin-wise faithfulness trends for each method in Figure 7.
 As shown in Table 6, BooookScore achieves the lowest sensitivity (0.0), making it the closest to zero among the four methods.
-Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary.
+Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the final bin, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary.
 These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon.
-Nevertheless, while BooookScore serves as a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation.
-This highlights the need for future work to develop methods that maintain faithfulness when generating long text directly from long contexts.
+In addition, we report results for other evaluation dimensions on the summaries generated using BooookScore in Table 7.
+The results show that BooookScore not only mitigates the Hallucinate at the Last phenomenon, but also improves the overall quality of the generated summaries.
+Nevertheless, while BooookScore provides a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation.
+This underscores the need for future methods that can maintain faithfulness when generating long text directly from long contexts.
 6 Related Work
@@
 (2023), decoding Shi et al.
-(2024); Wang et al.
+(2024b); Wang et al.
 (2025), or prompting Zhou et al.
 (2023).
-While effective in reducing overall hallucination rates, these studies typically do not analyze the within-sequence distribution of errors.
-The concept of positional bias in generation errors has been observed in related domains.
-Notably, studies on Large Vision-Language Models (LVLMs) have indicated that hallucinations tend to increase towards the end of generated textual descriptions of images Liu et al.
+In addition, Wan et al.
+(2025) analyzed positional hallucinations in long-context summarization.
+While these approaches reduce overall hallucination rates, most exisiting studies do not examine how hallucinations are distributed within the generated sequence, especially in long-form outputs.
+Positional bias in generation errors has also been observed in related domains.
+In particular, studies on Large Vision–Language Models (LVLMs) report that hallucinations tend to increase toward the end of generated image descriptions Liu et al.
 (2024a); Lee et al.
@@
 Our work presents the first dedicated study of where hallucinations occur in LLM-based long document summarization, moving beyond detection to uncover positional patterns, underlying causes, and mitigation strategies.
-Figure 7: Faithfulness scores for long summaries generated by different mitigation methods on the Wikipedia dataset across increasing output lengths.
+Methods Faithfulness(%) Completeness(%) Conciseness(%)
+Llama3.1-8B 90.33 100.0 41.35
+BooookScore 94.33 100.0 47.83
+Table 7: Overall faithfulness, completeness and conciseness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
 7 Conclusion
@@
 We investigated the contributing factors and explored targeted mitigation strategies.
-Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into positionally-aware generation techniques.
+Our work highlighted the importance of the output’s faithfulness and motivates future research into positionally-aware generation techniques.
 Limitations
@@
 Despite these limitations, we believe this work offers an important foundation for future research in the relatively underexplored domain of long-form output generation, particularly in summarization.
+Ethics Statement
+This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs.
+All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results.
+We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary.
+By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
 Acknowledgement
 This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
-Appendix A Details for Generating Summaries
-Building upon Section 2, this section details our approach to generating long summaries and presents an analysis of the generated outputs.
+Appendix A Attention Weight Calculation
+This section provides a more detailed and formal explanation of our attention weight computation method introduced in Section 4.
+Let the full sequence, consisting of the input context and the generated output summary, be defined as:
+$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
+where $L$ is the number of tokens.
+We partition $x_{full}$ into $K$ non-overlapping blocks of 100 tokens each, such that each block is represented as:
+$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
+for $k\in\{1,2,...,K\}$.
+Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
+$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
+$x^{middle}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
+$x^{last}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
+where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
+We define the average attention weight from block $x^{k}$ to each of the three sentences as follows:
+$\begin{gathered}\textrm{{Attn}}_{first}(k)=\frac{1}{100\cdot{T_{1}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{1}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{first})\\ \textrm{{Attn}}_{middle}(k)=\frac{1}{100\cdot{T_{2}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{2}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{middle})\\ \textrm{{Attn}}_{last}(k)=\frac{1}{100\cdot{T_{3}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{3}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{last})\end{gathered}$ (7)
+Appendix B Details for Atomic Fact Generation
+This section details our experimental details and human evaluation results from Section 2.
 Prompt
-Standard outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range.
-The specific word ranges used for prompting long-summary generation are provided in Table 7, and the prompt template employed for summary generation is shown in Table 10.
-Context Length Words Range
-1K 200 to 250
-2K 400 to 500
-3K 600 to 750
-4K 800 to 1000
-5K 1000 to 1250
-6K 1200 to 1500
-7K 1400 to 1750
-8K 1600 to 2000
-Table 7: Words range used in the prompt for long output generation.
-Generated Summary Length
-Our experiments revealed that controlling the output length of LLMs remains a significant challenge.
-This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger.
-In particular, with longer contexts, LLMs frequently generate long summaries that contain repeated sentences or even entire paragraphs.
-To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
-Table 8 presents the average word counts of both standard and long summaries generated by each model.
-As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
-Models Avg. Words
-standard long
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 8: The average word counts of both standard and long summaries generated by each model on the Wikipedia 7K dataset.
-Appendix B Details for Model Usage
+The prompt template employed for atomic fact generation is shown in Table 14.
+Human Evaluation
+Furthermore, we assess the quality of the atomic facts generated by the LLM through human evaluation, as reported in Table 8, focusing on atomic facts extracted from long summaries generated by Llama.
+For each context length, we randomly sampled 20 summaries, and the authors manually conducted the evaluation.
+Each atomic fact was independently annotated by the authors as either true or false, depending on whether it was supported by the corresponding summary.
+In total, 1,569 atomic facts were evaluated, and the two annotators agreed on 1,541 of them, corresponding to a 98.2% raw agreement.
+Because the task involves binary labels with a skewed class distribution, chance-corrected agreement coefficients are known to become unstable or difficult to interpret.
+We therefore report raw agreement as a clearer and more reliable indicator of annotation consistency in this setting.
+Context Length Validation(%)
+4K 95.92
+5K 99.01
+6K 96.18
+7K 99.12
+8K 99.64
+Table 8: Human evaluation of the quality of atomic facts.
+Appendix C Details for Model Usage
 In this section, we present the experimental details for the model usage in Section 3.
@@
 All summaries were generated using greedy decoding with float16 precision.
-Appendix C More Results with Different Methods
+Details for Bins
+Throughout all experiments, we used five bins.
+The bins were computed using the cut function in the pandas package, and the scores of the atomic facts were assigned to bins using open interval boundaries, meaning that the boundary values were not included in the bins.
+To ensure that our findings are not the result of the specific binning choice, we additionally report results using a finer-grained setting with ten bins in Table 10.
+The same trend persists even with this increased resolution, confirming that the Hallucinate at the Last phenomenon is robust to the choice of bin granularity.
+Models Generated Summary Bins Sensitivity
+1 2 3 4 5 6 7 8 9 10
+Llama3.1-8B 0.80 0.79 0.83 0.77 0.71 0.74 0.72 0.83 0.76 0.70 7.22
+Qwen2.5-7B 0.85 0.85 0.80 0.82 0.85 0.82 0.89 0.81 0.89 0.86 -1.78
+Table 10: Faithfulness scores and sensitivity for long summaries generated by Llama and Qwen.
+Appendix D More Results with Different Methods
 We adopt a diverse set of methods to validate the robustness of our faithfulness evaluation framework.
@@
 In contrast, the Qwen model does not yield the highest error rates in the final bin for both the cases of standard and long summaries.
-Appendix D More Results on Varying Datasets
-We report comparison of faithfulness scores in standard summaries generated by the Llama model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Figure 8.
+Appendix E More Results on Varying Datasets
 Minicheck Based
@@
 As demonstrated by the experimental results, consistent with the main results, none of the summaries generated by the Llama3.1-8B-Instruct model across various domains exhibited a sensitivity value below zero.
-Appendix E Attention Weight Calculation
-This section provides a more detailed and formal explanation of our attention weight computation method introduced in Section 4.
-Let the full sequence, consisting of the input context and the generated output summary, be defined as:
-$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
-where $L$ is the number of tokens.
-We partition $x_{full}$ into $K$ non-overlapping blocks of 100 tokens each, such that each block is represented as:
-$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
-for $k\in\{1,2,...,K\}$.
-Let the output summary contain three target sentences: the first, middle, and last sentence, denoted respectively as:
-$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
-$x^{second}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
-$x^{third}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
-where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
-We define the average attention weight from block $x^{k}$ to each of the three sentences as follows:
-$\begin{gathered}\textrm{{Attn}}_{first}(k)=\frac{1}{100\cdot{T_{1}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{1}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{first})\\ \textrm{{Attn}}_{middle}(k)=\frac{1}{100\cdot{T_{2}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{2}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{middle})\\ \textrm{{Attn}}_{last}(k)=\frac{1}{100\cdot{T_{3}}}\sum_{i=1}^{100}\sum_{j=1}^{T_{3}}\textrm{{attn}}(x_{i}^{k}\to x_{j}^{last})\end{gathered}$ (7)
 where $\textrm{{attn}}(x_{i}\to x_{j})$ denotes the attention weight from token $x_{i}$ to token $x_{j}$ when generating the output.
@@
 For LongWriter, we employed the LongWriter-Llama3.1-8B model [^8] https://huggingface.co/THUDM/LongWriter-llama3.1-8b with greedy decoding and bfloat16 precision.
-Appendix G Use of Large Language Model
-This paper primarily utilized the AI assistant (ChatGPT-5) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
+Appendix G Details for Generating Summaries
+Building upon Section 2, this section details our approach to generating long summaries and presents an analysis of the generated outputs.
+Prompt
+Standard outputs were standardized to a length of 100 to 200 words, while long outputs were approximated by converting 30% of the total context token length into a corresponding word range.
+The specific word ranges used for prompting long-summary generation are provided in Table 11, and the prompt template employed for summary generation is shown in Table 13.
+Context Length Words Range
+4K 800 to 1000
+5K 1000 to 1250
+6K 1200 to 1500
+7K 1400 to 1750
+8K 1600 to 2000
+Table 11: Words range used in the prompt for long output generation.
+Generated Summary Length
+Our experiments revealed that controlling the output length of LLMs remains a significant challenge.
+This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger.
+In particular, with longer contexts, LLMs often produce excessively long summaries containing repeated sentences or even entire paragraphs.
+To ensure fair and reliable evaluation, we excluded such outputs and considered only summaries with comparable lengths in our analysis.
+Specifically, we compared summaries within each sentence-count range and selected those with the highest sentence counts as length-matched summaries.
+For example, in our setting, the length-matched summaries correspond to 28 sentences for the 4K-context case and 40 sentences for the 8K-context case.
+Table 12 presents the average word counts of both standard and long summaries generated by each model.
+As the results indicate, LongWriter-Llama3.1-8B frequently produces considerably longer outputs, while Qwen2.5-7B-Instruct tends to generate relatively shorter summaries.
+Models Avg. Words
+standard long
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 12: The average word counts of both standard and long summaries generated by each model on the Wikipedia dataset.
+Appendix H Use of Large Language Model
+This paper primarily utilized the AI assistant (ChatGPT-5.1) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
 Input Prompt
@@
 Table 13: Prompt template used to generate summaries from the original document in Section 2.
-Figure 8: Comparison of faithfulness scores in standard summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.
+Input Prompt
+Decompose the text into atomic facts. {{Text}}
+Table 14: Prompt template used to generate atomic facts from the summary in Section 2.
 Models Context Generated Summary Bins Sensitivity